### v1

# NUDT4MSTAR: A New Dataset and Benchmark Towards SAR Target Recognition in the Wild

Abstract
Synthetic Aperture Radar (SAR) stands as an indispensable sensor for Earth observation, owing to its unique capability for all-day imaging. Nevertheless, in a data-driven era, the scarcity of large-scale datasets poses a significant bottleneck to advancing SAR automatic target recognition (ATR) technology. This paper introduces NUDT4MSTAR, a large-scale SAR dataset for vehicle target recognition in the wild, including 40 target types and a wide array of imaging conditions across 5 different scenes. NUDT4MSTAR represents a significant leap forward in dataset scale, containing over 190,000 images—tenfold the size of its predecessors. To enhance the utility of this dataset, we meticulously annotate each image with detailed target information and imaging conditions. We also provide data in both processed magnitude images and original complex formats. Then, we construct a comprehensive benchmark consisting of 7 experiments with 15 recognition methods focusing on the stable and effective ATR issues. Besides, we conduct transfer learning experiments utilizing various models trained on NUDT4MSTAR and applied to three other target datasets, thereby demonstrating its substantial potential to the broader field of ground objects ATR. Finally, we discuss this dataset’s application value and ATR’s significant challenges. To the best of our knowledge, this work marks the first-ever endeavor to create a large-scale dataset benchmark for fine-grained SAR recognition in the wild, featuring an extensive collection of exhaustively annotated vehicle images. We expect that the open source of NUDT4MSTAR will facilitate the development of SAR ATR and attract a wider community of researchers.
Index Terms:
Synthetic aperture radar (SAR), automatic target recognition (ATR), classification, detection, dataset, benchmark
1 Introduction
Can humans perceive and recognize objects on the Earth beyond the limitations of optical vision using microwave vision [20]? Synthetic aperture radar (SAR) [72, 61], an active imaging sensor, acquires the ground scattering characteristics by emitting the electromagnetic microwave. It has emerged as an indispensable tool for information acquisition in Earth observation and target surveillance [22, 76, 20], as its resolution remains unaffected by the observation distance or lighting conditions. Moreover, it has the advantage of operating effectively in various weather conditions. SAR interpretation has witnessed remarkable progress, propelled by the advancement of high-resolution imaging and artificial intelligence technologies [20, 64]. SAR automatic target recognition (ATR), which autonomously detects and classifies objects of interest (e.g., vehicles, ships, aircraft, and buildings) within imagery, is a critical frontier in advancing the capabilities of SAR interpretation. It has various civil and military applications such as urban management, disaster assessment, emergency rescue, and global survenlience [44, 18, 49, 25, 98]. Since the 1970s, researchers have been dedicated to developing a generalized stable SAR ATR system. However, the recognition of objects from SAR imagery have encountered significant challenges.
Figure 1: Motivation of our NUDT4MSTAR. Subfigure (a) shows the number of articles on SAR ATR in the last decade. As the first SAR target classification dataset, MSTAR has been the mainstream data benchmark. Subfigure (b) illustrates the data diversity of NUDT4MSTAR compared to MSTAR and its potential to become a new data benchmark.
Figure 2: Timeline of SAR classification dataset. Compared to other datasets focusing on target slices in simple scenes, we provide diverse target samples from different scenes to study the more difficult issues, and it is a much larger SAR ATR dataset.
Special characteristics. SAR imagery has special electromagnetic scattering attributions but is devoid of common features such as color and texture typically present in human vision. The scattering properties of targets are highly influenced by a multitude of factors, including the imaging angle, resolution, polarization, background, and the geometry structure and material of targets themselves [67, 39]. Additionally, SAR images suffer quality issues, such as speckle noise, geometric distortion, and image Defocus. These factors make it difficult to interpret SAR data, a recognition system independent of human vision.
Insufficient samples. Acquiring SAR images encompassing many categories under diverse operating conditions presents a significant challenge, and annotating SAR images is considerably more costly and labor-intensive than RGB images. Three decades have elapsed since the advent of the first SAR classification dataset in the 1990s, the Moving and Stationary Target Acquisition and Recognition (MSTAR) [2] in Fig. 2. The largest publicly accessible SAR target classification dataset remains limited to a small set of no more than 30,000 images. Besides, recent target classification datasets (OpenSARShip [30] and FUSAR-Ship [29]) have severe sample imbalance, ranging from a few to thousands of samples per class. A substantial amount of work remains to be undertaken for SAR target datasets.
Nonstandard benchmarks. A standardized dataset and experimental settings serve as the north star for developing advanced SAR ATR techniques. However, the mainstream [^1] For counting SAR ATR papers, we use the search formula “SAR target recognition (Title) or SAR ATR (Title) or SAR target classification (Title) or SAR classification (Title) and radar (Topic) ” on Web of Science. We added “and MSTAR (Topic)” to the previous search formula to count the number related to MSTAR. MSTAR dataset in Fig. 1 has dozens of distinct experimental settings, and existing algorithms have nearly reached near-saturation performance on this ideal benchmark [39]. Other classification datasets, such as OpenSARShip and FUSAR-Ship, also exhibit different experimental settings and dataset splitting. Moreover, the limiting access to non-open-source dataset settings and methods further intensifies this challenge.
Researchers have struggled to construct a large number of SAR target datasets to study generalized stable SAR ATR algorithms. Over the past few years, there has been a remarkable surge in remote sensing datasets across various modalities, including visible light [4, 51], SAR [52, 85], infrared [90], and multimodal [58]. As depicted in Fig. 2, recent SAR classification datasets have significantly improved data diversity. In particular, SAR detection datasets [52, 85] with 100,000 images have emerged. Yet, regarding SAR target recognition, no dataset has emerged that can completely surpass the MSTAR benchmark dataset, released in the 1990s, regarding the diversity of target types, imaging conditions, and data formats it offers. Notably, the lack of open-source availability, inadequate samples, long-tailed distributions, unabundant collection conditions, and independent experimental settings are still all issues that warrant attention. We suggest that a large-scale fine-grained SAR target dataset and benchmarks to investigate the electromagnetic scattering properties of targets across diverse conditions represent a critical and enduring endeavor.
Figure 3: NUDT4MSTAR. We collect target samples from 21 classes and 40 types of targets in 5 scenes with 3 modalities of sensors. Detailed target information is illustrated in Fig. 7. Due to the limitation of the scene size, there are two settings for target deployment, the first with 40 types in city, factory, and sandstone; the second with 11 types in woodland and bare soil. These vehicle types cover most current civilian vehicles in structure, size, and usage. In difficult scenarios (city and factory), vehicle positions are adjusted during acquisition to avoid pseudo-correlation between the target and the background. Although we focus on target characteristics in the SAR sensor, we also collect data from other sensors for future analysis of their differences.
As a first step towards building a large-scale SAR target dataset, we introduce NUDT4MSTAR, a fine-grained dataset for SAR vehicle target recognition in Fig. 3. In particular, NUDT4MSTAR collects 194,324 object images in the wild from 40 target types, 5 scenes, and various imaging conditions with detailed annotation [^2] This dataset focuses on the fine-grained single target recognition problem in SAR ATR, i.e., the image classification task, but we provide a horizontal bounding box label for the detection task.. It is the largest public SAR vehicle recognition dataset, even 10 times the size of any similar vehicle dataset previously available in Fig. 2. Then, to facilitate innovation and comparisons, we build a well-designed benchmark with 7 experimental settings and 15 recognition methods. Besides, the transfer learning results of various models trained on this dataset to other datasets show that our fine-grained recognition task benefits the feature learning of different ground targets. Finally, We provide detailed analyses and discussions to explore important issues for future research.
Our work shows the huge challenges and potential in SAR recognition and provides the benchmark for stable fine-grained ATR algorithms. The dataset and benchmark are available on our GitHub at https://github.com/waterdisappear/NUDT4MSTAR. We believe that our new dataset, with comprehensive imaging conditions and experiment benchmark, will further promote the promise and reproducible research in SAR ATR. The main contributions of this paper are three points:
1. We build the largest dataset for SAR vehicle target recognition in microwave vision. With various target types and imaging conditions, NUDT4MSTAR not only provides a systematic study of the effects of different conditions on SAR target characteristics but also promotes the development of new ATR technologies and research topics.
2. We constructed a target classification system for this vehicle dataset, including 4 categories, 21 classes, and 40 types. A transfer learning experiment based on pre-training on this dataset shows that a large fine-grained task can provide good weight initializations for recognizing ground targets.
3. We build a comprehensive benchmark with 7 experimental settings and 15 recognition methods. This benchmark provides detailed analysis and discussion that can guide SAR recognition algorithms’ designs and facilitate reproducible comparisons.
The remainder of this paper is organized as follows. Sec. 2 discusses related work in SAR target classification datasets and algorithms. Sec. 3 introduces our dataset NUDT4MSTAR. Secs. 4 conduct extensive experiments to build the benchmark. Sec. 5 concludes the paper and discusses future work.
2 Related Work
Here, we review the datasets and algorithms for SAR target classification and finally discuss the significance of the datasets for algorithmic studies.
TABLE I: Statistics of the SAR target classification datasets from the 1990s to 2020s. We focus on the electromagnetic scattering properties of individual objects, i.e., the most relevant to the classification dataset and task. Existing SAR target classification datasets face challenges such as no open source, insufficient samples, long-tailed distributions, and single collection conditions. OA.: Open access. # Classes.: Number of object classes. # Types.: Number of object types. Img. Size: image size. # Img.: Number of images. Res.: Resolution. Pol.: Polarization. “-” refers to unknown.
Dataset Year OA Source Band Pol. # Classes # Types Res. (m) Img. Size # Img. Description
[…]
CV Domes [16] 2010 $\checkmark$ simulation X quad 3 10 0.3 - - 3D simulation civil vehicle
Gotcha [15] 2012 $\checkmark$ airborne X - 7 13 0.3 - - 3D civil vehicle dataset
SARSIM [59] 2017 $\checkmark$ simulation X - 7 14 0.1 139 21,168 Simulation CAD vehicle
[…]
MGTD [5] 2019 - laboratory X single 1 2 0.01 128 1,728 Scaled vehicle models
SAMPLE [42] 2019 $\checkmark$ simulation X single 7 10 0.3 128 2,690 Simulation and measured pairs
[…]
IRIS-SAR [1] 2020 - simulation - - 8 355 - 512 63,900 Simulation CAD models
NUAAminiSAR [101] 2022 - airborne X - 9 9 0.1 - - Multi-angle circular SAR vehicle
MATD [81] 2022 $\checkmark$ airborne Ku - 2 2 - 128 144 Multi-angle strimap SAR aircraft
[…]
2.1 Datasets for SAR target classification
Datasets are the cornerstone of developing new techniques for SAR ATR, especially in a deep learning-based era. In the past decades, the MSTAR [2] dataset has played a key role in SAR target classification tasks. However, current deep learning-based methods are saturated on this dataset benchmark setting. MSTAR’s insufficient data volume also suffers from data biases such as background correlation [39, 47]. Therefore, many datasets have been built to advance the development of this field, especially in recent years. In order to provide a comprehensive analysis, we discuss SAR target classification datasets that focus on individual target characteristics from the 1990s to the 2020s, which are detailed listed in Table I.
Vehicle dataset. Vehicle datasets usually include various types of fine-grained vehicles and imaging conditions to study the robustness classification under different distributions in the training and test sets. We summarize the target classes and types of these SAR vehicle datasets in Fig. 5. MSTAR was released by the Defense Advanced Research Projects Agency (DARPA) as the first public dataset for SAR target classification. It contains SAR image slices of 10 military vehicles and a reference target under different imaging angles with spotlight SAR. However, MSTAR has a background correlation problem and ideal collection conditions. Compared to MSTAR, the QinetiQ [68] dataset provided a non-idealized acquisition condition with a non-centered target location and stronger background clutter. CV Domes [16] and Gotcha [15] constructed simulation and measurement samples of civilian vehicles. These two datasets are in the raw echo data format and can be used for SAR imaging research, but they are not as convenient for researchers to use as MSTAR. However, these two civilian vehicle datasets are not rich in target categories and acquisition conditions. SARSIM [59] and SAMPLE [42] are two simulation datasets used to study transfer learning. In particular, SARSIM uses 14 CAD models, and SAMPLE has simulated and measured sample pairs based on partially MSTAR’s public vehicle. The simulation data have discrepancies with the measured data in background clutter, target scattering characteristics, and speckle noise. Although MSTAR includes 3 grass locations, most public images are from the same location. MGTD [5] collects the training and test sets in two different laboratory backgrounds. Another issue is that the depression angles of MSTAR public images are mostly in 15° and 17°, and some categories of images are missing at other depression angles, such as 30° and 45°. Hence, NUAAminiSAR [101] collects images of military targets at different depression angles by circular SAR in a national defense park, including 8 military vehicles and an airplane. As shown above, no vehicle datasets are sufficiently rich in target types, scene settings, and sensor conditions.
Ship dataset. SAR ship datasets have emerged with the development of satellite SAR technology. The difference with the vehicle dataset is that the ship classification dataset usually uses a satellite platform, and the imaging angles are limited. Ship datasets can be constructed and labeled with global ship tracking systems, but label noise is a difficult problem. OpenSARShip [30], HR4S [3], and FUSAR-Ship [29] advance the development of ship classification methods. These datasets are derived from different satellites and face difficult harbor backgrounds and class imbalance challenges. OpenSARShip collects medium-resolution images of 16 classes in harbors with Sentinel-1. HR4S constructs high-resolution ship images in ports and coastal areas with GF-3 and RADARSAT-2. FUSAR-Ship contains high-resolution images under various sea backgrounds with 15 ship main classes and 98 subclasses based on GF-3. However, OpenSARShip and FUSAR-Ship have a serious class imbalance and long-tailed distribution challenges. Researchers construct various experimental settings and use some target categories for the classification based on these two datasets.
Aircraft dataset. Aircraft datasets have emerged in recent years, presenting unique challenges in remote sensing. Due to their streamlined surfaces and less obvious scattering characteristics, aircraft are more difficult to classify than vehicle and ship targets. Multiangle Aircraft Target Dataset (MATD) [81] collects SAR images of two aircraft at different imaging angles with a drone strimap SAR equipment. SAR-ACD [73] contains 2 categories and 6 types of civil aircraft in 3 airports.
Others. IRIS-SAR [1] is rich in aircraft, ships, and vehicle simulation types. It considers noise, clutter, and shadows in SAR image simulation. However, it does not take into account occlusion in complicated backgrounds.
Object detection dataset. SAR has many object detection datasets besides the classification datasets used for individual targets. These datasets are used for detecting multiple targets and have large image sizes. They contain more background interference and intensive targeting challenges. SAR object detection datasets similar to SARDet-100K [52] and FAIR-CSAR [85] with a volume of about 100,000 images are mainly satellite-based datasets, with a resolution of mostly greater than 1 m. This resolution makes it hard to detect and classify vehicle targets. Their non-cooperation objectives make it difficult to label the target types and accurately control the characteristics of the target under different angles and scenes. Therefore, our dataset provides a more fine-grained SAR vehicle characterization dataset under target class and imaging conditions.
Due to the high collection and annotation cost, collecting and publishing target samples under different imaging conditions is very difficult and time-consuming. Even after close to thirty years since its inception, MSTAR remains the most used classification dataset. The evolution of SAR ATR necessitates a new large-scale dataset and benchmark in the new era. After the above investigation, we suggest that a good SAR ATR dataset should pay attention to the following key points.
  1. $\bullet$
    Richness. The dataset should include rich target categories, scenes, and imaging conditions to approximate the real world and reduce data bias.
  2. $\bullet$
    Metadata. The dataset needs enough metadata, such as complex data, scene description, and imaging parameters. Researchers can propose new algorithms by combining metadata with SAR image characteristics.
  3. $\bullet$
    Standardization. An official benchmark and setting are beneficial to fair comparisons across different methods. In addition, the dataset format should be compatible with mainstream datasets for expansion and scaling.
  4. $\bullet$
    Usability. The dataset should be readily accessible and user-friendly for researchers to download and utilize. Some datasets may be difficult to access or require additional SAR imaging and pre-processing steps.
Figure 4: Importance of Dataset. Taking the MSTAR dataset as an example, subfigures (a1)-(a3) describe the common data inputs of existing algorithms. The MSTAR dataset provides processed magnitude data, raw complex data, and corresponding metadata. Therefore, this dataset remains the most used one over two decades after its release. Subfigures (b1)-(b2) show its mainstream experimental settings that are based on various imaging conditions of the MSTAR dataset. Other experimental settings include target configuration/version variation and simulation settings (simulation noise erosion and occlusion) [50]. However, the limited number of MSTAR categories and acquisition conditions restrict the further study of the SAR ATR. For example, the current occlusion usually uses zero values to fill some image regions, but the real occlusion in our measured dataset subfigure (b3) is not the same as the existing simulation. Therefore, building a new large-scale fine-grained vehicle dataset in the 2020s is necessary for the development of SAR target recognition based on the strengths of the previous datasets. (pseudo-color for better visualization of SAR magnitude images, and subfigures (b) are cropped for better visualization of target signature variations.)
Therefore, we build and open-source a new large-scale dataset, NUDT4MSTAR, with various categories and imaging conditions. This dataset is carefully labeled with extensive metadata and has an official benchmark based on 7 experimental settings and 15 comparison methods with 2 data formats. This dataset is our first step in building a new SAR target dataset system. In the future, we will expand the objects from vehicles to a wider category list to develop SAR ATR.
2.2 Review of SAR Target Classification Algorithms
Early target classification relies on hand-crafting features and classifiers, while contemporary methodologies are predominantly anchored in deep learning-based methods. According to their different feature representations, we divided target recognition methods into four distinct paradigms: traditional features, deep learning, feature fusion, and classification with metadata. In an era increasingly dominated by data-driven algorithms, the availability of a rich, standardized, and usable dataset enriched with metadata is becoming increasingly pivotal.
Traditional features. Previous studies have proposed a plethora of valuable features, encompassing geometric structure features [33, 23], electromagnetic-scattering features [63, 37, 11, 93, 45], time-frequency features [71], local descriptors [13, 70], and sparse representations [28]. These features are particularly tailored to the special coherent imaging mechanisms. For instance, SAR-HOG [70] employs a ratio-based gradient computation rather than the traditional differential gradient approach, and Attributed Scattering Centers (ASC) provides a feature set including location, geometry, and polarization information. Notably, extracting electromagnetic-scattering features relies on complex images with phase information, underscoring the necessity of providing comprehensive data formats.
Deep learning. Deep learning algorithms excel at uncovering intricate patterns within datasets, particularly when ample data is available. Nowadays, these algorithms achieve a remarkable nearly 99% accuracy on MSTAR’s basic Standard Operating Condition (SOC) settings, but the Extended Operating Condition (EOC) with robustness and few-shot are more challenging problems. Convolutional Neural Network (CNN) [40, 57] is a common network in SAR target classification. All-convolutional network (A-ConvNet) [8] was proposed to reduce overfitting and had high accuracy under many experiment settings but is not robust to noise. In recent years, Transformer [78, 14] has also been used in SAR target classification. After a decade of significant progress, SAR target recognition now incorporates many advanced deep learning techniques such as data augmentation [12, 41, 24], attention mechanism [65, 80], domain alignment [50], reinforcement Learning [35, 7], transfer learning [89], meta-learning [21], physical deep learning [9, 82, 32, 53], semi-supervised learning [97], and self-supervised learning [92, 34, 48]. The majority of research endeavors are based on the MSTAR dataset, and a subset of studies also create the different non-open source settings of OpenSARShip and FUSAR-Ship datasets. Nonetheless, the reliance on outdated datasets and the lack of standardized benchmarks are impeding the advancement of SAR target recognition in the 2020s.
Feature fusion. In order to improve the robustness and interpretability of deep learning methods, many feature fusion methods combining traditional and deep features have been proposed. Kechagias et al. [38] discussed the fusion of deep Learning and sparse coding. FEC [94] used deep CNN and ASC features based on complex data with amplitude and phase. Li et al. [43] combined scattering centers (SC) with graph convolutional network. Besides feature fusion methods, ESF [19] applied ASC features to guide deep model training. EFTL [54] proposed using the ASC model as weight initialization of a complex convolutional network. PIHA [31] proposed a hybrid attention mechanism with ASCs and the deep learning framework. LDSF [88] constructed a heterogeneous graph for the first time to fully exploit both the scattering information of the target components and their interactions. However, many SAR target datasets only provide magnitude images, ignoring phase information and quantization issues.
Classification with metadata. Imaging angles and target-related metadata are the two most prevalent categories of ancillary data. MSTAR and SAMPLE datasets furnish the imaging angles for each respective image. RotANet [84] combined predicting the rotational pattern of an image sequence with self-supervised learning. Zhang et al. [95] investigated the problem of optimal image imaging angle selection in few-shot recognition based on the MSTAR dataset. AaDRL [35] explored the SAR active recognition framework to fully exploit target characteristic variations at different azimuth angles based on the SAMPLE dataset. In addition, many multi-view methods [62, 74] were proposed with image sequences from different imaging angles. Since the target types are known, the researchers used 3D models and shape parameters as metadata for fine-grained tasks. For example, SARbake [60] used 3D CAD models and projection relations for semantic segmentation, and DBAE [83] utilized size and structure attributes for zero-shot classification. Consequently, a SAR target dataset enriched with pertinent metadata, encompassing both imaging parameters and detailed target characteristics, becomes an indispensable step in pursuing accurate and robust SAR target recognition.
Experimental settings. Indeed, the MSTAR dataset has more than 50 types of experimental configurations for SOC and EOC, as detailed in [39]. These configurations include image angle variation, target version/configure variation, noise corruption, occlusion, resolution variation, and background clutter variation [39, 50]. In a similar vein, other SAR classification datasets, such as OpenSARShip and FUSAR-Ship, lack a standard experimental benchmark, and researchers also devise different bespoke settings that leverage subsets of the target class for classification [32, 53].
Evaluation metrics. Classification accuracy and confusion matrix are the primary metrics employed for evaluating the performance of classification models. Accuracy quantifies the proportion of correctly classified samples relative to the total number of samples in the dataset. The confusion matrix describes the correspondence between the classification result and the ground truth. Within this matrix, each column corresponds to the predicted class labels, whereas each row represents the true class labels. Besides, the Receiver Operating Characteristic (ROC) curve is a valuable tool for assessing used for binary classification and detection tasks such as suppressing background clutter and outlier rejection [8].
Figure 5: Taxonomic systems of NUDT4MSTAR. For civilian vehicles, the taxonomy is based on Chinese and European vehicle classification standards [75, 17], according to the vehicle’s purpose, structure, size, and mass. For military vehicles, we followed the MSTAR taxonomic system [67]. (a) Taxonomic systems of NUDT4MSTAR. Our dataset is a comprehensive civilian vehicle map covering 4 main categories, 21 classes, and 40 types. We provide a detailed illustration of the histogram distribution for these 40 vehicle types, demonstrating the breadth and depth of our dataset. Its sufficient and balanced samples make it well-equipped to meet various experimental settings and studies. (b) Target classes of SAR vehicle datasets. We statistically analyze the number of civilian and military vehicle classes and types in SAR vehicle datasets. Our dataset greatly enhances the civilian vehicle richness for the SAR ATR, compared to the measured dataset Gotcha and the simulated datasets CVDomes and SARSim. (c) List of vehicle abbreviations.
We noticed a pronounced preference for the MSTAR dataset among SAR classification algorithms. Most algorithms are based on the MSTAR dataset, with a small number using other datasets such as SAMPLE, OpenSARShip, and FUSAR-Ship. This preference is attributed to the distinctive benefits of the MSTAR dataset released in the 1990s. As depicted in Fig. 4 (a), studies requiring complex data have opted for the MSTAR dataset, whereas other datasets that only provide magnitude data cannot be used to validate their algorithms. In addition, information about vehicle types in the dataset is also available on the web. Compared to other datasets with long-tailed distributions, the MSTAR dataset has a more balanced target sample and diverse imaging conditions, which facilitates the establishment of experimental frameworks and the investigation of recognition challenges. Therefore, MSTAR has accumulated a large number of algorithms and experimental settings. However, this dataset has not released all target samples, and the extant algorithms have achieved peak performance in prevalent settings, such as classification with ten target types, imaging angle variation, and version variation. The ideal acquisition conditions of the MSTAR dataset cannot show that the recognition problem is truly well solved. Scholars are endeavoring to address these limitations by proposing novel experimental settings and datasets. However, as shown in Fig. 4 (b), some simulation experimental settings based on measured data may not accurately reflect the real world. Besides, the sample diversity is difficult to ensure due to collection and labeling costs. Therefore, we perceive a constraint on advancing SAR ATR techniques due to data limitations. Constructing a new open-source dataset benchmark with diverse target categories and imaging conditions is the cornerstone of excellent and stable SAR ATR techniques.
3 NUDT4MSTAR
TABLE II: Statistics of NUDT4MSTAR. We segment the targets with fixed-size slices and random offsets to investigate their characteristics and aim for robust recognition under different collection conditions. We varied the target location in urban and factory scenes and collected data at the factory several times. In addition, some targets can not be labeled due to heavy occlusion. # Types.: Number of object types. Res.: Resolution. Pol.: Polarization. Dep.: Depression angle. Azi.: Target azimuth angle interval. Img. Size: image size. # Img.: Number of images in ground and slant range coordinate systems.
Scene # Types Platform Mode Band Res. (m) Pol. Dep. (°) Azi. (°) Img. Size # Img.
[…]
The emergence of new datasets and papers has led to rapid progress in various areas of SAR ATR, demonstrating the attention given to this issue in recent years. However, a large-scale dataset with various target categories and imaging conditions is still lacking because of SAR’s expensive costs and annotation. We aim to achieve as much diversity as possible in target categories, scene settings, and sensor conditions. With this in mind, our dataset has huge potential to advance the field of SAR ATR significantly. Its values for SAR ATR are:
  1. $\bullet$
    To provide a challenging and rich dataset for SAR vehicle recognition;
  2. $\bullet$
    To systematically study the effects of different conditions on SAR target characteristics;
  3. $\bullet$
    To give a benchmark to facilitate method innovation and comprehensive comparisons;
  4. $\bullet$
    To promote the development of new ATR technologies and research topics.
In this section, we describe the dataset construct program and then provide the details on NUDT4MSTAR in terms of key aspects, including data acquisition, professional system, statistical analysis, and dataset value.
3.1 Construct program
We aim to extend the dataset diversity as much as possible regarding targets, scenes, and sensors. These aspects are derived from the MSTAR’s discussion [67] on SAR target characteristics.
Figure 6: Illustrations of data acquisition. We annotate and cut to build target slices with corresponding metadata information. The range dimension of slant range complex data is in the line of sight direction, which results in the deformation of the target shape in this dimension. Therefore, we also provide ground range images after ground projecting based on slant range amplitude data.
Target. According to the Chinese and European vehicle classification standards [75, 17], First, we confirm the main target categories based on uses and structures of vehicles, including people carriers (car and bus), goods carriers (truck), and special purpose vehicles (special). Then, we subdivide each category into various target classes based on size and special structure and collect as different types as possible. Finally, NUDT4MSTAR has 4 categories, 21 classes, and 40 types. It covers the most of civilian vehicle categories and has the most richness on civil target classes compared to other SAR vehicle datasets in Fig. 5.
Figure 7: Demo images, classes, types, and sizes of forty vehicles in NUDT4MSTAR. Here are the SAR (pseudo-color for better visualization of SAR magnitude images) and corresponding RGB images. As we can see, the obvious difference between the vehicles in the SAR images is the scattering characteristic variation due to size and structure. Therefore, we classify various vehicle types based on their structure and size. Their length, width, and height measurements are listed in meters (m). Besides, it is worth noting that our SAR images are not in the ideal situation where the target is located right in the center of the demo image, just as in the MSTAR dataset, but has a random offset similar to the QinetiQ dataset.
Scene. Most previous target datasets were collected in open scenes such as grasslands, focusing on target characteristics and lacking significant background interference. Therefore, we discuss both difficult and simple scenes [^3] Since we mainly consider single target characteristics, discussing densely distributed targets such as car parks are not included in this version. in Fig. 10. The urban and factory have varying occlusion degrees and other strong scattering targets, such as buildings and roadside trees. Dense trees in woodlands can weaken target scattering or increase the intensity of non-target scattering. Bare soil and sandstone are clean backgrounds, although sandstone occasionally has strong scattering background points. These interferences have a stronger effect on the target features and are more compatible with real-world challenging recognition tasks.
Sensor. The common sensor conditions are variations in depression and azimuth angle, and we collect target samples at different angles as shown in Fig. 11. We provide a balanced sample under different depression angles, including all target types, compared to MSTAR’s several targets. However, our azimuth angles are not as dense as MSTAR in most scenes due to the limitation of strimap imaging. In addition, we provide data in different polarizations and bands to study their effects on SAR target characteristics. We simultaneously gather data from satellites and airborne to study the platforms and resolutions.
Table II shows the open-source data list. After a year and a half’s effort (from April 2023 to October 2024), we completed program designation, data acquisition and processing, and benchmark construction. We commissioned the Institute of Stealthy Engineering Technology Company in Beijing to complete the data collection and annotation.
3.2 Data acquisition
Here, we present our data acquisition pipeline in terms of imaging, annotation, and product.
Imaging. We use the Unmanned Aerial Vehicle (UAV) platform to carry sensor equipment and collect data. Two antennas acquire quad polarization radar echo in the X and Ku bands. Besides, the Position and Orientation System (POS) device provides the Global Positioning System (GPS) and Inertial Measurement Unit (IMU) information for motion compensation. After imaging processing, we have slant range complex images. Ground range data is obtained by ground projection based on slant range amplitude images and POS information.
Annotation. We annotate the target types and positions with rectangular box labels according to the optical images and placement positions. Since we focus on individual target characteristics, the targets are placed at a certain distance from each other. Besides our vehicles, other vehicles in the scene are labeled as “other”. After labeling, we acquired the target slices and added a random offset.
Product. We offer data products in two coordinate systems. The distance dimension of the slant range data is the line of sight direction, and we provide original complex data. The ground range images is projected to the ground truth distance and is processed with nonlinear quantization. The corresponding annotation files include basic image information and target, scene, and sensor parameters.
3.3 Professional system
We construct a target taxonomic system to facilitate subsequent extensions and integration with other datasets. Data format and metadata are also discussed, considering the specificity of SAR imaging regarding view, steps, and operating conditions.
Figure 8: Distribution of object size. Our dataset has a richer object sizes compared to another civilian vehicle dataset (Gotcha). It is clear that the size and ratio of the vehicles are within a certain range due to real-world constraints, such as official regulations and driving requirements, and we will continue to expand the object categories beyond vehicles.
[…]
Metadata system. Considering the effects of target, scene, and sensor on SAR images, we add many factors to the annotation to facilitate various recognition methods with metadata. Image information includes basic image sizes and formats. The target attributes include taxonomic systems, target size, and image position. Although we discuss single target characteristics with classification tasks, rectangular boxes also support detection tasks. Sensor parameters have the depression angle, target azimuth angle, band, polarization, and resolution in two dimensions. However, we hope that researchers consider the difficulty and accuracy of obtaining metadata when making full use of them. For example, there is the potential for angular confusion when measuring target azimuths and the potential for vehicle size to change after modifications.
3.4 Statistical Analysis
Figure 9: Non-centered target slices. We apply the non-centered target setting following the QinetiQ dataset to increase detection difficulty.
Figure 10: Influence of different scenes on target characteristics in SAR images. Most current SAR target classification datasets are collected in simple scenes, e.g., (a) MSTAR is collected in a flat grass scene with a clean background and pseudo-correlation between target and background. While (b) our NUDT4MSTAR samples are collected under different positions in various scenes. In Fig. (b), it is clear that the same target under the same imaging angle has obvious signature changes due to different scenes. For example, other objects higher up in front of and behind the target can create occlusions and layovers, reducing reflected energy and increasing non-target scattering. Besides, the target shadow in difficult scenes is not as obvious as the MSTAR data. (c) Occlusion and layover. Compared to the bare soil scene without interfering objects, the factory and woodland scenarios indicate occlusion and layover. Shadows from interfering objects in the former may obscure targets, while trees in the latter are likely to increase non-target scattering. Due to occlusion and layover are complex results of a combination of target, interference, and imaging geometry, we illustrate this problem with a single target demo. These statistics are from an SUV Chang’an CSCS75 Plus (Vehicle 12) at different image angles in the 3 scenes’s ground range images.
Figure 11: Target azimuth angle distribution of (a) MSTAR and (b) NUDT4MSTAR. Compared to the MSTAR dataset, which has depression angles mainly publicized at 17° and 15° with incomplete angles for most target classes, our NUDT4MSTAR dataset provides balance and comprehensive angles for all targets. However, the sampling interval of target angles is sparser due to strip imaging.
Figure 12: Ratio of strong and weak points in different bands and polarization. It can be noticed that the band and polarization can have a statistically significant effect on the target signatures. These statistical data are the target region inside the vehicle’s rectangular box of sandstone scenes. Points with pixel values greater than or equal to 128 are treated as strong scattering points for 8-bit ground range images.
Figure 13: Complex data. Here is a data example of Vehicle 5 (Great Wall Voleex C50).
We now discuss the properties of our dataset and provide some statistics compared with others.
Class distribution. NUDT4MSTAR includes 4 categories, 21 classes and 40 types of vehicles with balanced samples, as shown in Fig. 5. Relative to extant SAR target datasets, our dataset’s targets’ enhanced diversity and intricacy introduce novel formidable challenges to the SAR target fine-grained recognition research. Moreover, adequate samples guarantee the availability of a substantial array of target classes across experimental settings, facilitating more diverse investigations.
Object size. The length, width, and height of our 40 objects are listed in Fig. 7, and we can see that they have various structures, different sizes, and similar size ratios in Fig. 8. Compared to another civilian vehicle dataset (Gotcha), our dataset has a wider range of target classes and size distribution.
Reference target. Besides deploying corner reflectors for resolution measurements, we produce and place a reference target of multiple geometries as the “SLICY” target in the MSTAR dataset for scattering issues of a standard geometry.
Other vehicle targets. Other vehicle targets in the scene are also labeled for the potential detection task in large remote images. However, this dataset mainly discusses the scattering properties of individual targets. These interfering vehicles can be used as part of a future rejection issue.
Non-centered location. Most SAR target datasets place the target in the image center or contain only the target region, but the QinetiQ dataset uses a non-centered setting. Since remote sensing differs from the human eye view, which customarily centers the object of interest, the overhead view requires searching targets with more interference. As shown in Fig. 9, we randomly added position offsets to increase the recognition difficulty with detecting non-centered target locations.
Occlusion in scenes. Previous vehicle datasets focus on target characteristics with simple scenes, and we collect samples from different scenes with the same target and angle. Fig. 10 shows their different influences with occlusions and layovers in different scene regions. The shadow of buildings and roadside trees in the factory may obscure targets and reduce target scattering, and the nearby trees in woodland have more obvious occlusion and increase non-target scattering.
Imaging angle. Different imaging angles affect the target scattering characteristics, and the imaging geometry relationship can change the interference in the scene. Our dataset has various imaging angles with balance distribution in different depression angles, as shown in Fig. 11. However, the sampling interval of target azimuth angles is not equally dense in different scenes due to the limitation of stripmap imaging and cost problems.
Band and polarization. Table I shows most vehicle datasets with the MSTAR dataset’s X-band and HH polarization settings. However, Fig. 12 demonstrates that this setting is not the only one to consider, as band and polarization have a noticeable effect on target scattering. We provide two bands and quad polarization to support discussing these sensor parameters.
Complex data. Feature extraction based on complex data is a hot research concern, so we provide complex data in slant distance coordinates stored in .mat format as Fig. 13 shown.
Number of image bits. RGB images are stored in 8-bit format, whereas SAR images have a larger range of digital data values. Therefore, we provide different bit formats to investigate quantification for enhancing weaker target scattering points.
Cross-platform problem. We collect a small amount of satellite data synchronously under the same scenes and targets to investigate the difficulty of recognition across platforms.
Quality control. To ensure high-quality data, we considered challenges such as target and site hire, airspace applications, weather conditions, and data checking. Eight people completed the data acquisition over six months. In addition, 14 labelers and two inspectors carried out the data annotation over about four months. The entire project cost about CNY 1.45 million.
3.5 Dataset value
Based on the diversity of target categories, scenes, imaging conditions, and data format with detailed annotation, we recommend that this dataset be used for SAR target recognition studies in the following areas.
Robustness recognition. Our data has diverse acquisition conditions, and each image is labeled with specific information. Therefore, we can discuss in detail the robustness of the SAR recognition algorithms when the acquisition conditions (e.g., scene, and sensor) vary between the training and test sets. If we consider data from different acquisition conditions as domains, the training set can include multiple domains to learn domain invariant features, i.e., the domain generalization [99].
Few-shot learning. In addition to acquisition condition variations, the problem of small samples due to high costs is also a concern issue. The few-shot setting can be combined with robust recognition, and our dataset’s large number of target types can enrich the task construction of meta-learning [27, 87, 86] settings in existing few-shot learning.
Transfer learning. Our dataset contains a large diversity of samples that can be used to study the transfer problem of the pre-training model with different SAR target datasets. In addition, the large number of samples can further increase the volume of SAR target samples to advance the study of self-supervised learning [56, 48] and foundation models [6, 46].
Incremental learning. The diversity of our dataset can support the study of domain incremental learning and class incremental learning [77]. The domain incremental learning can improve the algorithm’s robustness with a dynamic process, and class incremental learning can progressively increase the ability to recognize or reject new classes. This incremental capability is required for SAR ATR systems facing open environments.
Physical deep learning. SAR images have unique properties, such as complex phase and polarization. Our dataset provides a variety of data formats that can support recognition studies based on these unique characteristics rather than relying only on quantified SAR magnitude images. In addition, detailed metadata can also provide more gains for SAR ATR.
We encourage researchers to propose new experimental settings and research issues based on this dataset. Don’t hesitate to contact us if you get new ideas.
4 Benchmarks
TABLE III: SOC and EOC settings of NUDT4MSTAR. The imaging conditions in SOC are similar, while EOC considers variations in a single imaging condition. Simple scenes are sandstone and bare soil, and complex scenes are urban, factory, and woodland. The separate labeling of the ground distance and slant distance images results in their numbers of annotations not being strictly corresponding. # Types.: Number of object types. Dep.: Depression angle. Azi.: Target azimuth angle. Pol.: Polarization. # Img. (Ground): Number of images in ground range. # Img. (Slant): Number of images in slant range.
Setting Set # Types Scene Dep. Azi. Band Pol. # Img. (Ground) # Img. (Slant)
[…]
We consider 7 experimental settings with 2 data formats as classification and detection benchmarks for the NUDT4MSTAR dataset in Table III. These experimental settings include 2 SOC settings sampled from similar distributions and 5 EOC settings with obvious distribution shifts. The data formats are magnitude images in the ground range coordinate system and complex images in the slant range coordinate system.
In conclusion, we evaluate our benchmarks under 7 experimental settings and 15 recognition methods with classification and detection tasks based on magnitude images and classification tasks based on complex images. These experiments provide valuable insights and huge potential into recognizing SAR target characteristics under different imaging conditions and fine-grained classes. Our dataset is sufficient to give a new challenging benchmark to advance SAR target recognition technology. We also welcome researchers to contact us to construct richer experimental benchmarks from original data and publish them with your clear insights.
4.1 Experimental settings
Here, we describe the experimental settings in Table III, including 2 SOC settings sampled from similar distributions and 5 EOC settings with obvious condition and distribution shifts.
[…]
Compared methods. We select classical or recent open-source work in computer vision and SAR target recognition, in addition to recently proposed methods from our laboratory. For classification, we have chosen 6 computer vision methods (VGG16 [69], ResNet18 [26], ResNet34 [26], ConvNeXt [57], ViT [14], and HiViT [96]) and 4 SAR ATR methods (HDANet [50], SARATR-X [46, 48], MS-CVNets [91], and LDSF [88]). MS-CVNets and LDSF are complex images methods others are magnitude methods. VGG16, ResNet18, ResNet34, ConvNeXt, ViT, HiViT, and SARATR-X all use the pre-training weights. For detection, we have chosen 4 computer vision methods (Faster RCNN [66], CenterNet [100], YOLOv8 [36], and YOLOv10 [79]) and 2 SAR ATR methods (DiffDet4SAR [98], and SARATR-X [46, 48]). These methods are based on magnitude images. Because of the long-term dataset construction process, we will keep updating the dataset homepage with new experimental settings and method results.
Transfer learning. To show that this dataset can improve the model’s performance for SAR target feature extraction, we perform training on SOC-40 according to Table IV, and the weights are used for other SAR target classification datasets as initialization for fine-tuning. Fine-tuning parameter settings follow our previous paper [48], and the baseline is the weight of ImageNet supervised pre-training for these models.
4.2 Results and analyses
Here, we present the results of our classification, detection, and transfer learning tasks.
TABLE IV: Classification results. We use accuracy as a metric, i.e., the number of correctly classified samples in proportion to the total number of samples. Gray color methods use complex images, and magnitude image methods are sorted by network architecture, such as CNN and Transformer. Bolded text indicates the best result, while underlined text is the next best result.
Setting VGG16 [69] HDANet [50] ResNet18 [26] ResNet34 [26] ConvNeXt [57] ViT [14] HiViT [96] SARATR-X [46, 48] MS-CVNet [91] LDSF [88]
SOC-40 88.8 89.1 90.6 91.7 96.0 76.4 86.8 96.4 80.9 88.2
SOC-50 72.9 63.7 71.2 72.9 81.6 59.2 68.0 85.2 55.2 65.7
EOC-Scene 21.6 33.6 16.1 18.0 16.5 12.9 15.8 19.5 26.1 29.7
– (city) 22.7 33.7 18.1 18.9 17.9 11.6 13.9 20.4 26.1 32.7
– (factory) 20.5 34.3 13.7 16.8 14.4 15.3 17.8 18.4 25.6 30.1
– (woodland) 6.8 27.6 4.2 5.4 4.2 1.97 3.8 2.7 20.2 26.3
EOC-Depression 33.2 32.9 33.9 37.2 43.1 30.4 31.4 39.9 21.8 28.3
– (30) 52.4 54.0 52.2 55.9 63.2 43.7 47.3 58.1 39.2 46.5
– (45) 33.5 32.7 34.2 38.3 45.1 31.0 32.0 41.2 19.2 25.9
– (60) 14.0 12.3 15.4 17.4 21.3 16.7 15.0 20.5 8.1 12.7
EOC-Azimuth 15.7 16.4 14.9 16.5 21.1 29.0 22.8 26.4 19.0 22.5
– (60) 20.7 19.4 18.0 22.1 26.2 34.4 27.2 28.9 26.8 31.0
– (120) 8.2 9.5 8.3 8.5 13.5 26.4 14.2 22.8 13.2 15.6
– (180) 17.6 17.9 16.4 17.4 19.9 20.8 22.2 23.0 16.9 19.7
– (240) 7.11 7.0 7.4 7.9 10.6 20.3 11.1 12.3 10.9 12.6
– (300) 26.0 30.4 25.5 27.5 36.6 44.4 40.9 46.7 28.7 33.7
EOC-Band 78.8 79.5 83.1 83.8 88.4 65.7 70.7 89.2 63.7 60.9
– inverse 74.5 79.8 76.1 79.0 84.6 62.3 78.4 89.1 60.9 64.2
EOC-Polarization 72.5 63.1 71.4 70.5 83.1 53.6 67.1 84.6 49.0 55.1
– VV 72.4 63.0 72.9 72.1 84.4 55.2 69.1 87.5 52.0 63.4
– HV 72.3 62.7 70.2 69.4 82.4 52.4 65.9 83.1 45.8 49.0
– VH 72.8 63.5 71.0 70.0 82.7 55.1 66.5 83.3 46.1 52.8
Figure 14: Results under SOC-50 with different training set ratio. We find a linear trend decreasing test set accuracy as training decreases on a logarithmic scale.
Classification. From different methods in Table IV, we note that deep learning methods can achieve high performance in SAR fine-grained classification, such as the recent methods ConvNeXt and SARATR-X in SOC-40. Methods based on complex images, such as MS-CVNet and LDSF, better distinguish between targets and clutter in EOC-Scene. The model recognition performance and robustness of the latest 2020s CNN architecture ConvNeXt are superior to the previous CNN models. ConvNeXt is more robust under EOC-Deprssion than other methods, and ViT is insensitive to azimuthal angle change. Our previously proposed method, HDANet, improves the attention mechanism and uses a segmentation task so it can be used in different scenarios. In addition, our proposed LDSF introduces a scattering graph structure that effectively improves the discrimination of targets and clutter. SARATR-X is a pre-trained self-supervised learning model on SAR with the HiViT architecture, and its performance is significantly improved compared to HiViT. Besides, the Transformer models are trained with some low-layer weights frozen, as we found that the full update tends to lead to overfitting.
The different experimental settings in Table IV show that there are still great challenges in SAR target recognition. As the number of fine-grained classes increased, previous methods were no longer able to achieve nearly 99% accuracy on MSTAR’s SOC. SOC-50 is closer to the actual situation, as we are unlikely to obtain a large number of target samples under different conditions. The test set accuracy tends to decrease linearly as the training samples are reduced on a logarithmic scale in Fig 14. In the difficult scene of EOC-Scene, deep learning encounters great challenges, especially in the background clutter interference inside the trees inside the woodland. The significant deterioration in accuracy indicates that deep learning models trained on simple backgrounds have difficulty distinguishing between targets and complex background clutter. The use of complex information or the inclusion of auxiliary tasks such as segmentation during training can enhance the recognition ability, but the alteration of target features due to background clutter interference is still a difficult challenge. We observe a significant decrease using the complex data compared to the magnitude data under angles and sensor parameter variations. These results are because the complex phase is more sensitive to these parameters’ transformations, and the magnitude of the data is additionally preprocessed to enhance the image’s visual quality. Changes in imaging angle have a greater impact on target characteristics than band and polarization, especially large depression and azimuth angle changes. Different imaging angles lead to variations in target and background clutter, as well as different occlusions and scattering resulting from changes in imaging geometry. We have not found a good method to deal with the above scene and angle changes. The shift in target features due to band and polarization can be solved very effectively by ConvNeXt and SARATR-X. However, due to cost issues, we have not discussed the effect of different platform resolutions on recognition in the current version, which is another important issue.
TABLE V: Detection results. We use mAP50 [55] as a metric, i.e., the average precision at IoU 0.5 because the horizontal bounding box does not exactly fit the contours of the different targets. Bolded text indicates the best result, while underlined text is the next best result.
Setting CenterNet [100] YOLOv8 [36] YOLOv10 [79] Faster RCNN [66] DiffDet4SAR [98] SARATR-X [46, 48]
SOC-40 83.5 86.1 91.7 88.3 90.9 95.5
SOC-50 74.7 64.4 63.5 65.0 76.1 79.6
EOC-Scene 30.2 26.2 21.5 35.5 37.2 20.4
– (city) 37.8 30.8 24.8 42.6 45.9 18.1
– (factory) 27.1 24.3 20.3 32.9 35.2 24.1
– (woodland) 23.8 25.1 20.0 30.7 32.8 25.2
EOC-Depression 33.5 22.7 22.3 29.0 31.0 33.3
– (30) 57.3 42.9 42.5 50.1 52.1 55.4
– (45) 34.2 21.4 21.6 27.7 30.1 35.0
– (60) 8.1 4.4 4.1 6.4 8.2 9.9
EOC-Azimuth 24.5 18.1 19.2 26.9 28.7 22.9
– (60) 21.4 15.9 18.6 26.8 28.9 20.5
– (120) 20.0 13.0 14.3 20.8 22.1 22.3
– (180) 18.7 15.0 14.1 19.0 20.1 18.6
– (240) 14.9 10.4 11.4 18.0 19.0 14.0
– (300) 56.0 44.6 44.7 57.9 60.1 46.7
EOC-Band 69.1 72.4 74.5 72.9 75.1 87.5
– inverse 75.9 75.2 72.1 69.5 73.8 83.7
EOC-Polarization 74.8 56.0 59.2 58.3 61.2 75.3
– VV 75.6 56.9 62.4 57.9 60.4 77.4
– HV 74.3 55.3 57.6 59.0 62.1 73.9
– VH 75.1 56.4 58.3 58.5 61.0 74.8
Detection. As illustrated in Table V, we find that the two-stage approach (Faster RCNN, DiffDet4SAR, and SARATR-X) has better robustness than the single-stage approach (CenterNet, YOLOv8, and YOLOv10) under most EOC conditions. In particular, CenterNet models the target center using a Gaussian kernel, demonstrating superior detection capabilities under EOC-Depression and EOC-Polarization conditions. However, YOLO series methods, which rely on fixed grid partitioning and coordinate regression, exhibit limitations in detection performance under EOC conditions. In contrast, our proposed DiffDet4SAR, a diffusion model-based detection framework, consistently outperforms other algorithms in EOC-SCENE and EOC-Azimuth scenarios. This advantage stems from the robust denoising box structure, which enhances target location accuracy, and the use of central difference convolution to amplify target saliency. The foundation model SARATR-X, based on self-supervised learning with SAR images, performs well for fine-grained recognition and most of the EOC conditions, but the performance for scene and angle changes needs to be improved.
Compared with the classification task, the detection task can effectively distinguish between target and background clutter, especially for target discovery in complex scenes such as woodlands. However, most of the detection performance is slightly lower than the classification performance due to the SAR image quality issues that result in noisy labels and rectangular box offsets. Besides, our fine-grained detection tasks, encompassing both accurate localization and precise classification, are inherently more complex than single precise classification tasks, necessitating advanced model architectures and sophisticated algorithm designs. Furthermore, although we changed the target position, such adjustments are not performed frequently due to acquisition time and cost constraints. Similarly, there is still much scope for improvement in robustness under scene and imaging angle variations.
Figure 15: Transfer learning results. We report the 5-shot fine-tuning performance of transfer learning on three SAR target classification datasets after NUDT4MSTAR training. The results show that our dataset has a good effect on ground target datasets such as vehicles and aircraft, but the effect is insignificant for sea targets.
Transfer learning. Fig 15 shows how our dataset can help different models benefit from feature learning and be used for target recognition in other datasets. We use Table’s SOC-40 trained model weights as initializations on other target datasets with 5-shot sample experiments [48]. The fine-tuning results show that similar ground target-type datasets, such as vehicles and airplanes, would greatly benefit from training on our data. However, the transfer effect is not obvious for the ship target class dataset in the sea scene.
4.3 Discussion
Researchers have pursued realizing a stable and efficient SAR ATR recognition system, but previous vehicle benchmark datasets have not provided difficult challenges. Our results show that current state-of-the-art methods still face many robustness and small sample difficulties. Robust feature extraction under different imaging conditions and scenes is still a worthy research problem. It is also important to efficiently use a small number of labeled samples and mine complex and polarization information. Moreover, there is no single method that can handle different perturbation conditions, demonstrating the complexity of stable SAR ATR. Of course, building a large-scale dataset and advancing SAR ATR technology is a long-term process. In a data-driven approach dominant in the 2020s, a large-scale dataset will help to extract stabilizing features and be effective in many other datasets and tasks. Based on the results, we discuss several potential directions for the future development of ATR systems.
Robustness under different conditions. As Table IV and  V show, the variations in imaging conditions present a significant challenge for recognition. This variability in the domain distribution requires the extraction of stable invariant target features under different conditions.
Efficient utilization of samples. Due to the difficulty of acquiring and labeling SAR images, it is unlikely that the sample size will reach the same size as the visible light dataset. Therefore, it is important to improve learning efficiency by utilizing a small number of samples in Fig 14.
Incorporation with multi-dimensional information. Using magnitude images alone loses many SAR image attributes, such as complex information that can effectively distinguish between targets and clutter in Table IV. Therefore, auxiliary information such as complex, polarization, and angle can be introduced in SAR ATR, but increasing the dimensionality of the information may increase the instability and complexity of the features.
Fast incremental learning capability. Because of the open and dynamic world, real ATR systems face constantly changing imaging conditions and target types. They need to effectively recognize anomalies and update their capabilities in response to variations. This data provides a rich set of target types and imaging conditions that can be used to explore this possibility.
Subsequently, we plan to construct a different resolution version and paired polarization data. We will continue to expand the scenes and target classes by combining satellite data with a semi-automated and low-cost method.
5 Conclusion
A large-scale fine-grained SAR vehicle dataset and benchmark, named NUDT4MSTAR, has been established in this paper. It contains various target types, scene variations, and imaging conditions and is 10 times the size of previous datasets of the same type. Our benchmark of 7 experiment settings and 15 methods result show that stable and efficient SAR ATR recognition method are still challengeable. It will advance diverse issue explorations and SAR ATR techniques with large-scale data. In the future, we will continue to contribute to the SAR target datasets, and we welcome researchers to contact us to jointly promote the progress of the SAR ATR field.

### v2

@@
-NUDT4MSTAR: A New Dataset and Benchmark Towards SAR Target Recognition in the Wild
+NUDT4MSTAR: A Large Dataset and Benchmark Towards Remote Sensing Object Recognition in the Wild
 Abstract
-Synthetic Aperture Radar (SAR) stands as an indispensable sensor for Earth observation, owing to its unique capability for all-day imaging.
+As an indispensable sensor for Remote sensing, Synthetic Aperture Radar (SAR) has a unique capability for all-day imaging.
 Nevertheless, in a data-driven era, the scarcity of large-scale datasets poses a significant bottleneck to advancing SAR automatic target recognition (ATR) technology.
-This paper introduces NUDT4MSTAR, a large-scale SAR dataset for vehicle target recognition in the wild, including 40 target types and a wide array of imaging conditions across 5 different scenes.
+This paper introduces NUDT4MSTAR, a large-scale SAR dataset for remote sensing target recognition in the wild, including 40 vehicle target types and various imaging conditions across 5 realistic scenes.
 NUDT4MSTAR represents a significant leap forward in dataset scale, containing over 190,000 images—tenfold the size of its predecessors.
-To enhance the utility of this dataset, we meticulously annotate each image with detailed target information and imaging conditions.
-We also provide data in both processed magnitude images and original complex formats.
+We meticulously annotate each image with detailed target information and imaging conditions.
+Besides, data in both processed magnitude images and original complex formats are provided.
 Then, we construct a comprehensive benchmark consisting of 7 experiments with 15 recognition methods focusing on the stable and effective ATR issues.
-Besides, we conduct transfer learning experiments utilizing various models trained on NUDT4MSTAR and applied to three other target datasets, thereby demonstrating its substantial potential to the broader field of ground objects ATR.
+Besides, we conduct transfer learning experiments utilizing various models training on NUDT4MSTAR and apply them to three other target datasets, demonstrating its substantial potential for the broader field of ground objects ATR.
 Finally, we discuss this dataset’s application value and ATR’s significant challenges.
@@
 Index Terms:
-Synthetic aperture radar (SAR), automatic target recognition (ATR), classification, detection, dataset, benchmark
+Remote sensing, transformer, benchmark dataset, performance evaluation, synthetic aperture radar, automatic target recognition, image classification, object detection, foundation model, deep learning
 1 Introduction
-Can humans perceive and recognize objects on the Earth beyond the limitations of optical vision using microwave vision [20]?
+Figure 2: Motivation of our NUDT4MSTAR.
+Subfigure (a) depicts the most frequent keywords in 21,780 journal papers published in the remote sensing field (TGRS, J-STARS, GRSL, ISPRS P&R, and JAG) from 2020 to 2024.
+The size of each word is proportional to its frequency, highlighting that concepts such as synthetic aperture radar (SAR), image classification, and object detection [21, 10, 99] have garnered substantial attention.
+Subfigure (b) focuses on the number of publications related to SAR ATR over the past five years, a cross-area of the concepts highlighted in Subfigure (a).
+As the pioneering dataset for SAR target classification, MSTAR has long served as the predominant benchmark due to its unique data diversity and accumulated benchmarks.
+However, the lack of large-scale datasets has significantly limited the growth of this research field in recent years.
+Remote sensing imagery interpretation is crucial for understanding our Earth comprehensively via diverse multi-modal sensors [64, 27, 111].
+Recently, the advent of the big data era has propelled the evolution of pre-trained foundation models, which have been widely applied across various remote sensing downstream tasks involving different multi-modal data [30, 107, 87, 26, 3, 58].
+Nevertheless, not all modal tasks in remote sensing possess a large number of target samples to exploit the potential of big data and foundation models.
 Synthetic aperture radar (SAR) [78, 36], an active imaging sensor, acquires the ground scattering characteristics by emitting the electromagnetic microwave.
-It has emerged as an indispensable tool for information acquisition in Earth observation and target surveillance [22, 76, 20], as its resolution remains unaffected by the observation distance or lighting conditions.
-Moreover, it has the advantage of operating effectively in various weather conditions.
-SAR interpretation has witnessed remarkable progress, propelled by the advancement of high-resolution imaging and artificial intelligence technologies [20, 64].
-SAR automatic target recognition (ATR), which autonomously detects and classifies objects of interest (e.g., vehicles, ships, aircraft, and buildings) within imagery, is a critical frontier in advancing the capabilities of SAR interpretation.
-It has various civil and military applications such as urban management, disaster assessment, emergency rescue, and global survenlience [44, 18, 49, 25, 98].
+As shown in Fig 2 (a), it has emerged as an indispensable tool for information acquisition in Remote sensing and Earth observation [23, 83, 21], as its resolution remains unaffected by the observation distance or lighting conditions.
+SAR automatic target recognition (ATR), which autonomously detects and classifies objects of interest (e.g., , vehicles, ships, aircraft, and buildings) within imagery, is a critical increasing frontier in SAR interpretation, as illustrated in Fig.
+2 (b).
+It has various civil and military applications, such as urban management, disaster assessment, emergency rescue, and global surveillance [10, 54, 28, 108].
+Can humans perceive and recognize objects on the Earth beyond the limitations of optical vision using microwave vision [21, 36, 49]?
 Since the 1970s, researchers have been dedicated to developing a generalized stable SAR ATR system.
-However, the recognition of objects from SAR imagery have encountered significant challenges.
-Figure 1: Motivation of our NUDT4MSTAR.
-Subfigure (a) shows the number of articles on SAR ATR in the last decade.
-As the first SAR target classification dataset, MSTAR has been the mainstream data benchmark.
-Subfigure (b) illustrates the data diversity of NUDT4MSTAR compared to MSTAR and its potential to become a new data benchmark.
-Figure 2: Timeline of SAR classification dataset.
-Compared to other datasets focusing on target slices in simple scenes, we provide diverse target samples from different scenes to study the more difficult issues, and it is a much larger SAR ATR dataset.
+However, recognizing objects within SAR images has encountered significant challenges in a big data era, especially the scarcity of large-scale target samples.
 Special characteristics.
 SAR imagery has special electromagnetic scattering attributions but is devoid of common features such as color and texture typically present in human vision.
-The scattering properties of targets are highly influenced by a multitude of factors, including the imaging angle, resolution, polarization, background, and the geometry structure and material of targets themselves [67, 39].
+The scattering properties of targets are highly influenced by a multitude of factors, including the imaging angle, resolution, polarization, background, and the geometry structure and material of targets themselves [73, 44, 69].
 Additionally, SAR images suffer quality issues, such as speckle noise, geometric distortion, and image Defocus.
-These factors make it difficult to interpret SAR data, a recognition system independent of human vision.
+These factors make it difficult to recognize objects in SAR, a unique recognition system independent of human vision.
 Insufficient samples.
 Acquiring SAR images encompassing many categories under diverse operating conditions presents a significant challenge, and annotating SAR images is considerably more costly and labor-intensive than RGB images.
-Three decades have elapsed since the advent of the first SAR classification dataset in the 1990s, the Moving and Stationary Target Acquisition and Recognition (MSTAR) [2] in Fig.
-2.
-The largest publicly accessible SAR target classification dataset remains limited to a small set of no more than 30,000 images.
-Besides, recent target classification datasets (OpenSARShip [30] and FUSAR-Ship [29]) have severe sample imbalance, ranging from a few to thousands of samples per class.
+Three decades have elapsed since the advent of the first SAR classification dataset in the 1990s, the Moving and Stationary Target Acquisition and Recognition (MSTAR) [2].
+As shown in Fig.
+3, the largest publicly accessible SAR target classification dataset remains limited to a small set of fewer than 30,000 images.
+Besides, recent classification datasets (OpenSARShip [34] and FUSAR-Ship [33]) have severe sample imbalance, ranging from a few to thousands of samples per class.
 A substantial amount of work remains to be undertaken for SAR target datasets.
@@
 A standardized dataset and experimental settings serve as the north star for developing advanced SAR ATR techniques.
-However, the mainstream [^1] For counting SAR ATR papers, we use the search formula “SAR target recognition (Title) or SAR ATR (Title) or SAR target classification (Title) or SAR classification (Title) and radar (Topic) ” on Web of Science.
-We added “and MSTAR (Topic)” to the previous search formula to count the number related to MSTAR.
-MSTAR dataset in Fig.
-1 has dozens of distinct experimental settings, and existing algorithms have nearly reached near-saturation performance on this ideal benchmark [39].
+However, the mainstream MSTAR dataset in Fig.
+2 (b) has dozens of distinct experimental settings, and existing algorithms have nearly reached near-saturation performance on this ideal benchmark [44].
 Other classification datasets, such as OpenSARShip and FUSAR-Ship, also exhibit different experimental settings and dataset splitting.
 Moreover, the limiting access to non-open-source dataset settings and methods further intensifies this challenge.
-Researchers have struggled to construct a large number of SAR target datasets to study generalized stable SAR ATR algorithms.
+Researchers have struggled to construct many SAR target datasets to study generalized stable SAR ATR algorithms.
 Over the past few years, there has been a remarkable surge in remote sensing datasets across various modalities, including visible light [5, 56], SAR [57, 93], infrared [98], and multimodal [65].
-As depicted in Fig.
-2, recent SAR classification datasets have significantly improved data diversity.
+Recent SAR classification datasets in Table II have significantly improved data diversity.
 In particular, SAR detection datasets [57, 93] with 100,000 images have emerged.
@@
 We suggest that a large-scale fine-grained SAR target dataset and benchmarks to investigate the electromagnetic scattering properties of targets across diverse conditions represent a critical and enduring endeavor.
-Figure 3: NUDT4MSTAR.
-We collect target samples from 21 classes and 40 types of targets in 5 scenes with 3 modalities of sensors.
-Detailed target information is illustrated in Fig.
-7.
-Due to the limitation of the scene size, there are two settings for target deployment, the first with 40 types in city, factory, and sandstone; the second with 11 types in woodland and bare soil.
-These vehicle types cover most current civilian vehicles in structure, size, and usage.
-In difficult scenarios (city and factory), vehicle positions are adjusted during acquisition to avoid pseudo-correlation between the target and the background.
-Although we focus on target characteristics in the SAR sensor, we also collect data from other sensors for future analysis of their differences.
-As a first step towards building a large-scale SAR target dataset, we introduce NUDT4MSTAR, a fine-grained dataset for SAR vehicle target recognition in Fig.
-3.
-In particular, NUDT4MSTAR collects 194,324 object images in the wild from 40 target types, 5 scenes, and various imaging conditions with detailed annotation [^2] This dataset focuses on the fine-grained single target recognition problem in SAR ATR, i.e., the image classification task, but we provide a horizontal bounding box label for the detection task..
-It is the largest public SAR vehicle recognition dataset, even 10 times the size of any similar vehicle dataset previously available in Fig.
-2.
+TABLE I: MSTAR v.s.
+NUDT4MSTAR.
+We can see the potential of the NUDT4MSTAR dataset as a new benchmark, surpassing MSTAR in terms of data diversity and challenges.
+AFRL: Air Force Research Laboratory.
+NUDT: National University of Defense Technology.
+# Types.: Number of object types.
+# Samples: Number of samples.
+Dataset MSTAR NUDT4MSTAR
+Year 1995 2025
+Country USA CHN
+Institution AFRL NUDT
+# Types 10 40
+Location ideally centered random, noncentered
+Scene grass 5 realistic scenes
+Platform Airborne Airborne
+[…]
+Band X X, Ku
+Polarization single quad
+Depression angle () 15, 17, 30, 45 15, 30, 45, 60
+[…]
+Annotation classification classification, detection
+# Samples 14,557 194,324
+As a first step towards building a large-scale SAR target dataset, we introduce NUDT4MSTAR [^1] The name NUDT4MSTAR comes from the abbreviation of the College of Electronic Science and Technology, the National University of Defense Technology, as Department IV.
+This dataset focuses on the finegrained single target recognition problem in SAR ATR, i.e., the image classification task, but we provide a horizontal bounding box label for the object detection task.
+NUDT4MSTAR: https://github.com/waterdisappear/NUDT4MSTAR., a fine-grained dataset for vehicle target recognition as a new breakthrough to the previous benchmark dataset MSTAR in Table.
+I.
+In particular, NUDT4MSTAR collects 194,324 object images in the wild from 40 target types, 5 scenes, and various imaging conditions with detailed annotation.
+It is the largest public SAR vehicle recognition dataset, even 10 times the size of any similar vehicle dataset previously available, as shown in Table.
+II.
 Then, to facilitate innovation and comparisons, we build a well-designed benchmark with 7 experimental settings and 15 recognition methods.
@@
 Our work shows the huge challenges and potential in SAR recognition and provides the benchmark for stable fine-grained ATR algorithms.
-The dataset and benchmark are available on our GitHub at https://github.com/waterdisappear/NUDT4MSTAR.
+The dataset and benchmark are available on our GitHub.
 We believe that our new dataset, with comprehensive imaging conditions and experiment benchmark, will further promote the promise and reproducible research in SAR ATR.
@@
 1.
-We build the largest dataset for SAR vehicle target recognition in microwave vision.
-With various target types and imaging conditions, NUDT4MSTAR not only provides a systematic study of the effects of different conditions on SAR target characteristics but also promotes the development of new ATR technologies and research topics.
+We build a large dataset for remote sensing object recognition in the Wild.
+NUDT4MSTAR includes 40 distinct vehicle target types and over 194,324 target samples across 5 realistic scenes and various imaging conditions.
+It provides a systematic study of the effects of different conditions on SAR target characteristics and promotes the development of new ATR research topics.
 2.
-We constructed a target classification system for this vehicle dataset, including 4 categories, 21 classes, and 40 types.
-A transfer learning experiment based on pre-training on this dataset shows that a large fine-grained task can provide good weight initializations for recognizing ground targets.
+We provide a comprehensive benchmark with 7 experimental settings and 15 recognition methods, covering both classification and detection recognition tasks.
+This benchmark offers a standardized experimental platform for researchers to compare the performance of different algorithms, facilitating innovation and development in SAR ATR technology.
 3.
-We build a comprehensive benchmark with 7 experimental settings and 15 recognition methods.
-This benchmark provides detailed analysis and discussion that can guide SAR recognition algorithms’ designs and facilitate reproducible comparisons.
+We demonstrate the NUDT4MSTAR dataset’s potential for foundation models.
+A transfer learning experiment based on various models pre-trained on this dataset shows that it is not only suitable for vehicle target recognition but also has the potential to support research in other related objects, promoting the application of SAR ATR in multiple situations.
 The remainder of this paper is organized as follows.
@@
 Here, we review the datasets and algorithms for SAR target classification and finally discuss the significance of the datasets for algorithmic studies.
-TABLE I: Statistics of the SAR target classification datasets from the 1990s to 2020s.
+2.1 Datasets for SAR target classification
+Datasets are the cornerstone of developing new techniques for SAR ATR, especially in a deep learning-based era.
+In the past decades, the MSTAR [2] dataset has played a key role in SAR target classification tasks.
+However, current deep learning-based methods are saturated on this dataset benchmark setting.
+MSTAR’s insufficient data volume also suffers from data biases such as background correlation [44, 52].
+Therefore, many datasets have been built to advance the development of this field, especially in recent years.
+In order to provide a comprehensive analysis, we discuss SAR target classification datasets that focus on individual target characteristics from the 1990s to the 2020s, which are detailed listed in Table II and Fig.
+3.
+TABLE II: Statistics of the SAR target classification datasets from the 1990s to 2020s.
 We focus on the electromagnetic scattering properties of individual objects, i.e., the most relevant to the classification dataset and task.
@@
 […]
-2.1 Datasets for SAR target classification
-Datasets are the cornerstone of developing new techniques for SAR ATR, especially in a deep learning-based era.
-In the past decades, the MSTAR [2] dataset has played a key role in SAR target classification tasks.
-However, current deep learning-based methods are saturated on this dataset benchmark setting.
-MSTAR’s insufficient data volume also suffers from data biases such as background correlation [39, 47].
-Therefore, many datasets have been built to advance the development of this field, especially in recent years.
-In order to provide a comprehensive analysis, we discuss SAR target classification datasets that focus on individual target characteristics from the 1990s to the 2020s, which are detailed listed in Table I.
+Figure 3: Timeline of SAR classification dataset.
+Compared to other datasets focusing on target slices in simple scenes, we provide diverse target samples from different scenes to study the more difficult issues, and it is a much larger SAR ATR dataset.
 Vehicle dataset.
@@
   1. $\bullet$
-    Richness. The dataset should include rich target categories, scenes, and imaging conditions to approximate the real world and reduce data bias.
-  2. $\bullet$
-    Metadata.
-The dataset needs enough metadata, such as complex data, scene description, and imaging parameters.
+    Richness.
+The dataset should include rich target categories, scenes, and imaging conditions to approximate the real world and reduce data bias.
+Besides, the dataset needs enough metadata, such as complex data, scene description, and imaging parameters.
 Researchers can propose new algorithms by combining metadata with SAR image characteristics.
@@
     Standardization.
-An official benchmark and setting are beneficial to fair comparisons across different methods.
-In addition, the dataset format should be compatible with mainstream datasets for expansion and scaling.
+An official dataset and benchmark are beneficial to fair research and comparisons.
+In addition, the dataset format and target system should be standardized to be compatible with mainstream datasets and for further expansion and scaling.
   3. $\bullet$
@@
 Traditional features.
-Previous studies have proposed a plethora of valuable features, encompassing geometric structure features [33, 23], electromagnetic-scattering features [63, 37, 11, 93, 45], time-frequency features [71], local descriptors [13, 70], and sparse representations [28].
+Previous studies have proposed a plethora of valuable features, encompassing geometric structure features [38, 24, 13], electromagnetic-scattering features [70, 42, 102, 50], time-frequency features [77], local descriptors [15, 76], and sparse representations [32].
 These features are particularly tailored to the special coherent imaging mechanisms.
@@
 In recent years, Transformer [85, 16] has also been used in SAR target classification.
-After a decade of significant progress, SAR target recognition now incorporates many advanced deep learning techniques such as data augmentation [12, 41, 24], attention mechanism [65, 80], domain alignment [50], reinforcement Learning [35, 7], transfer learning [89], meta-learning [21], physical deep learning [9, 82, 32, 53], semi-supervised learning [97], and self-supervised learning [92, 34, 48].
+After a decade of significant progress, SAR target recognition now incorporates many advanced deep learning techniques such as data augmentation [14, 46, 25], attention mechanism [71, 88], domain alignment [55], reinforcement Learning [40, 8], transfer learning [97], meta-learning [22], physical deep learning [11, 90, 37, 59], semi-supervised learning [106, 80], and self-supervised learning [101, 39, 53].
 The majority of research endeavors are based on the MSTAR dataset, and a subset of studies also create the different non-open source settings of OpenSARShip and FUSAR-Ship datasets.
@@
 Constructing a new open-source dataset benchmark with diverse target categories and imaging conditions is the cornerstone of excellent and stable SAR ATR techniques.
-3 NUDT4MSTAR
-TABLE II: Statistics of NUDT4MSTAR.
+3 Building NUDT4MSTAR: Why, What, How
+In this section, we present NUDT4MSTAR from three perspectives: (1) Why we constructed this dataset in the 2020s; (2) What we did to ensure its diversity and standardization; (3) How the diversity and statistical analysis of this dataset to support SAR ATR.
+TABLE III: Statistics of NUDT4MSTAR.
 We segment the targets with fixed-size slices and random offsets to investigate their characteristics and aim for robust recognition under different collection conditions.
@@
     To promote the development of new ATR technologies and research topics.
-In this section, we describe the dataset construct program and then provide the details on NUDT4MSTAR in terms of key aspects, including data acquisition, professional system, statistical analysis, and dataset value.
 3.1 Construct program
@@
 Target.
-According to the Chinese and European vehicle classification standards [75, 17], First, we confirm the main target categories based on uses and structures of vehicles, including people carriers (car and bus), goods carriers (truck), and special purpose vehicles (special).
-Then, we subdivide each category into various target classes based on size and special structure and collect as different types as possible.
+First of all, according to the Chinese and European vehicle classification standards [82, 19], we confirm the main target categories based on uses and structures of vehicles, including people carriers (car and bus), goods carriers (truck), and special purpose vehicles (special).
+Then, we subdivide each category into various target classes based on size and special structure and collect as many different types as possible.
 Finally, NUDT4MSTAR has 4 categories, 21 classes, and 40 types.
-It covers the most of civilian vehicle categories and has the most richness on civil target classes compared to other SAR vehicle datasets in Fig.
-5.
+It covers most civilian vehicle categories and has the most richness on civil target classes compared to other SAR vehicle datasets.
 Figure 7: Demo images, classes, types, and sizes of forty vehicles in NUDT4MSTAR.
@@
 We simultaneously gather data from satellites and airborne to study the platforms and resolutions.
-Table II shows the open-source data list.
+Table III shows the open-source data list.
 After a year and a half’s effort (from April 2023 to October 2024), we completed program designation, data acquisition and processing, and benchmark construction.
@@
 Data format and metadata are also discussed, considering the specificity of SAR imaging regarding view, steps, and operating conditions.
-Figure 8: Distribution of object size.
-Our dataset has a richer object sizes compared to another civilian vehicle dataset (Gotcha).
-It is clear that the size and ratio of the vehicles are within a certain range due to real-world constraints, such as official regulations and driving requirements, and we will continue to expand the object categories beyond vehicles.
 […]
@@
 3.4 Statistical Analysis
+We now discuss the properties of our dataset and provide some statistics compared with others.
+Class distribution.
+NUDT4MSTAR includes 4 categories, 21 classes and 40 types of vehicles with balanced samples, as shown in Fig.
+5.
+Relative to extant SAR target datasets, our dataset’s targets’ enhanced diversity and intricacy introduce novel formidable challenges to the SAR target fine-grained recognition research.
+Moreover, adequate samples guarantee the availability of a substantial array of target classes across experimental settings, facilitating more diverse investigations.
+Figure 8: Distribution of object size.
+Our dataset has a richer object size distribution compared to another civilian vehicle dataset (Gotcha).
+It is clear that the size and ratio of the vehicles are within a certain range due to real-world constraints, such as official regulations and driving requirements, and we will continue to expand the object categories beyond vehicles.
+Object size.
+The length, width, and height of our 40 objects are listed in Fig.
+7, and we can see that they have various structures, different sizes, and similar size ratios in Fig.
+8.
+Compared to another civilian vehicle dataset (Gotcha), our dataset has a wider range of target classes and size distribution.
+Reference target.
+Besides deploying corner reflectors for resolution measurements, we produce and place a reference target of multiple geometries as the “SLICY” target in the MSTAR dataset for scattering issues of a standard geometry.
+Other vehicle targets.
+Other vehicle targets in the scene are also labeled for the potential detection task in large remote images.
+However, this dataset mainly discusses the scattering properties of individual targets.
+These interfering vehicles can be used as part of a future rejection issue.
 Figure 9: Non-centered target slices. We apply the non-centered target setting following the QinetiQ dataset to increase detection difficulty.
+Non-centered location.
+Most SAR target datasets place the target in the image center or contain only the target region, but the QinetiQ dataset uses a non-centered setting.
+Since remote sensing differs from the human eye view, which customarily centers the object of interest, the overhead view requires searching targets with more interference.
+As shown in Fig.
+9, we randomly added position offsets to increase the recognition difficulty with detecting non-centered target locations.
 Figure 10: Influence of different scenes on target characteristics in SAR images.
@@
 These statistics are from an SUV Chang’an CSCS75 Plus (Vehicle 12) at different image angles in the 3 scenes’s ground range images.
-Figure 11: Target azimuth angle distribution of (a) MSTAR and (b) NUDT4MSTAR.
-Compared to the MSTAR dataset, which has depression angles mainly publicized at 17° and 15° with incomplete angles for most target classes, our NUDT4MSTAR dataset provides balance and comprehensive angles for all targets.
-However, the sampling interval of target angles is sparser due to strip imaging.
-Figure 12: Ratio of strong and weak points in different bands and polarization.
-It can be noticed that the band and polarization can have a statistically significant effect on the target signatures.
-These statistical data are the target region inside the vehicle’s rectangular box of sandstone scenes.
-Points with pixel values greater than or equal to 128 are treated as strong scattering points for 8-bit ground range images.
-Figure 13: Complex data. Here is a data example of Vehicle 5 (Great Wall Voleex C50).
-We now discuss the properties of our dataset and provide some statistics compared with others.
-Class distribution.
-NUDT4MSTAR includes 4 categories, 21 classes and 40 types of vehicles with balanced samples, as shown in Fig.
-5.
-Relative to extant SAR target datasets, our dataset’s targets’ enhanced diversity and intricacy introduce novel formidable challenges to the SAR target fine-grained recognition research.
-Moreover, adequate samples guarantee the availability of a substantial array of target classes across experimental settings, facilitating more diverse investigations.
-Object size.
-The length, width, and height of our 40 objects are listed in Fig.
-7, and we can see that they have various structures, different sizes, and similar size ratios in Fig.
-8.
-Compared to another civilian vehicle dataset (Gotcha), our dataset has a wider range of target classes and size distribution.
-Reference target.
-Besides deploying corner reflectors for resolution measurements, we produce and place a reference target of multiple geometries as the “SLICY” target in the MSTAR dataset for scattering issues of a standard geometry.
-Other vehicle targets.
-Other vehicle targets in the scene are also labeled for the potential detection task in large remote images.
-However, this dataset mainly discusses the scattering properties of individual targets.
-These interfering vehicles can be used as part of a future rejection issue.
-Non-centered location.
-Most SAR target datasets place the target in the image center or contain only the target region, but the QinetiQ dataset uses a non-centered setting.
-Since remote sensing differs from the human eye view, which customarily centers the object of interest, the overhead view requires searching targets with more interference.
-As shown in Fig.
-9, we randomly added position offsets to increase the recognition difficulty with detecting non-centered target locations.
 Occlusion in scenes.
@@
 The shadow of buildings and roadside trees in the factory may obscure targets and reduce target scattering, and the nearby trees in woodland have more obvious occlusion and increase non-target scattering.
+Figure 11: Target azimuth angle distribution of (a) MSTAR and (b) NUDT4MSTAR.
+Compared to the MSTAR dataset, which has depression angles mainly publicized at 17° and 15° with incomplete angles for most target classes, our NUDT4MSTAR dataset provides balance and comprehensive angles for all targets.
+However, the sampling interval of target angles is sparser due to strip imaging.
 Imaging angle.
@@
 However, the sampling interval of target azimuth angles is not equally dense in different scenes due to the limitation of stripmap imaging and cost problems.
+Figure 12: Ratio of strong and weak points in different bands and polarization.
+It can be noticed that the band and polarization can have a statistically significant effect on the target signatures.
+These statistical data are the target region inside the vehicle’s rectangular box of sandstone scenes.
+Points with pixel values greater than or equal to 128 are treated as strong scattering points for 8-bit ground range images.
 Band and polarization.
-Table I shows most vehicle datasets with the MSTAR dataset’s X-band and HH polarization settings.
+Table II shows most vehicle datasets with the MSTAR dataset’s X-band and HH polarization settings.
 However, Fig.
@@
 We provide two bands and quad polarization to support discussing these sensor parameters.
-Complex data. Feature extraction based on complex data is a hot research concern, so we provide complex data in slant distance coordinates stored in .mat format as Fig. 13 shown.
+Complex data. Feature extraction based on complex data is a hot research concern, so we provide complex data in slant distance coordinates stored in .mat format.
 Number of image bits.
@@
 Eight people completed the data acquisition over six months.
-In addition, 14 labelers and two inspectors carried out the data annotation over about four months.
-The entire project cost about CNY 1.45 million.
+In addition, 14 labelers and two inspectors perform the data annotation about four months.
 3.5 Dataset value
@@
 Our data has diverse acquisition conditions, and each image is labeled with specific information.
-Therefore, we can discuss in detail the robustness of the SAR recognition algorithms when the acquisition conditions (e.g., scene, and sensor) vary between the training and test sets.
+Therefore, we can discuss in detail the robustness of the SAR recognition algorithms when the acquisition conditions (e.g., scene and sensor sensors) vary between the training and test sets.
 If we consider data from different acquisition conditions as domains, the training set can include multiple domains to learn domain invariant features, i.e., the domain generalization [109].
@@
 4 Benchmarks
-TABLE III: SOC and EOC settings of NUDT4MSTAR.
+We consider 7 experimental settings with 2 data formats as classification and detection benchmarks for the NUDT4MSTAR dataset in Table IV.
+These experimental settings include 2 SOC settings sampled from similar distributions and 5 EOC settings with obvious distribution shifts.
+The data formats are magnitude images in the ground range coordinate system and complex images in the slant range coordinate system.
+In conclusion, we evaluate our benchmarks under 7 experimental settings and 15 recognition methods with classification and detection tasks based on magnitude images and classification tasks based on complex images.
+These experiments provide valuable insights and huge potential into recognizing SAR target characteristics under different imaging conditions and fine-grained classes.
+Our dataset is sufficient to give a new challenging benchmark to advance SAR target recognition technology.
+We also welcome researchers to contact us to construct richer experimental benchmarks from original data and publish them with your clear insights.
+4.1 Experimental settings
+Here, we describe the experimental settings in Table IV, including 2 SOC settings sampled from similar distributions and 5 EOC settings with obvious condition and distribution shifts.
+TABLE IV: SOC and EOC settings of NUDT4MSTAR.
 The imaging conditions in SOC are similar, while EOC considers variations in a single imaging condition.
@@
 […]
-We consider 7 experimental settings with 2 data formats as classification and detection benchmarks for the NUDT4MSTAR dataset in Table III.
-These experimental settings include 2 SOC settings sampled from similar distributions and 5 EOC settings with obvious distribution shifts.
-The data formats are magnitude images in the ground range coordinate system and complex images in the slant range coordinate system.
-In conclusion, we evaluate our benchmarks under 7 experimental settings and 15 recognition methods with classification and detection tasks based on magnitude images and classification tasks based on complex images.
-These experiments provide valuable insights and huge potential into recognizing SAR target characteristics under different imaging conditions and fine-grained classes.
-Our dataset is sufficient to give a new challenging benchmark to advance SAR target recognition technology.
-We also welcome researchers to contact us to construct richer experimental benchmarks from original data and publish them with your clear insights.
-4.1 Experimental settings
-Here, we describe the experimental settings in Table III, including 2 SOC settings sampled from similar distributions and 5 EOC settings with obvious condition and distribution shifts.
 […]
@@
 Transfer learning.
-To show that this dataset can improve the model’s performance for SAR target feature extraction, we perform training on SOC-40 according to Table IV, and the weights are used for other SAR target classification datasets as initialization for fine-tuning.
+To show that this dataset can improve the model’s performance for SAR target feature extraction, we perform training on SOC-40 according to Table V, and the weights are used for other SAR target classification datasets as initialization for fine-tuning.
 Fine-tuning parameter settings follow our previous paper [53], and the baseline is the weight of ImageNet supervised pre-training for these models.
@@
 Here, we present the results of our classification, detection, and transfer learning tasks.
-TABLE IV: Classification results.
+TABLE V: Classification results.
 We use accuracy as a metric, i.e., the number of correctly classified samples in proportion to the total number of samples.
@@
 – VH 72.8 63.5 71.0 70.0 82.7 55.1 66.5 83.3 46.1 52.8
-Figure 14: Results under SOC-50 with different training set ratio. We find a linear trend decreasing test set accuracy as training decreases on a logarithmic scale.
 Classification.
-From different methods in Table IV, we note that deep learning methods can achieve high performance in SAR fine-grained classification, such as the recent methods ConvNeXt and SARATR-X in SOC-40.
+From different methods in Table V, we note that deep learning methods can achieve high performance in SAR fine-grained classification, such as the recent methods ConvNeXt and SARATR-X in SOC-40.
 Methods based on complex images, such as MS-CVNet and LDSF, better distinguish between targets and clutter in EOC-Scene.
@@
 Besides, the Transformer models are trained with some low-layer weights frozen, as we found that the full update tends to lead to overfitting.
-The different experimental settings in Table IV show that there are still great challenges in SAR target recognition.
+The different experimental settings in Table V show that there are still great challenges in SAR target recognition.
 As the number of fine-grained classes increased, previous methods were no longer able to achieve nearly 99% accuracy on MSTAR’s SOC.
@@
 However, due to cost issues, we have not discussed the effect of different platform resolutions on recognition in the current version, which is another important issue.
-TABLE V: Detection results.
+Figure 13: Results under SOC-50 with different training set ratio. We find a linear trend decreasing test set accuracy as training decreases on a logarithmic scale.
+Detection.
+As illustrated in Table VI, we find that the two-stage approach (Faster RCNN, DiffDet4SAR, and SARATR-X) has better robustness than the single-stage approach (CenterNet, YOLOv8, and YOLOv10) under most EOC conditions.
+In particular, CenterNet models the target center using a Gaussian kernel, demonstrating superior detection capabilities under EOC-Depression and EOC-Polarization conditions.
+However, YOLO series methods, which rely on fixed grid partitioning and coordinate regression, exhibit limitations in detection performance under EOC conditions.
+In contrast, our proposed DiffDet4SAR, a diffusion model-based detection framework, consistently outperforms other algorithms in EOC-SCENE and EOC-Azimuth scenarios.
+This advantage stems from the robust denoising box structure, which enhances target location accuracy, and the use of central difference convolution to amplify target saliency.
+The foundation model SARATR-X, based on self-supervised learning with SAR images, performs well for fine-grained recognition and most of the EOC conditions, but the performance for scene and angle changes needs to be improved.
+TABLE VI: Detection results.
 We use mAP50 [61] as a metric, i.e., the average precision at IoU 0.5 because the horizontal bounding box does not exactly fit the contours of the different targets.
@@
 – VH 75.1 56.4 58.3 58.5 61.0 74.8
-Detection.
-As illustrated in Table V, we find that the two-stage approach (Faster RCNN, DiffDet4SAR, and SARATR-X) has better robustness than the single-stage approach (CenterNet, YOLOv8, and YOLOv10) under most EOC conditions.
-In particular, CenterNet models the target center using a Gaussian kernel, demonstrating superior detection capabilities under EOC-Depression and EOC-Polarization conditions.
-However, YOLO series methods, which rely on fixed grid partitioning and coordinate regression, exhibit limitations in detection performance under EOC conditions.
-In contrast, our proposed DiffDet4SAR, a diffusion model-based detection framework, consistently outperforms other algorithms in EOC-SCENE and EOC-Azimuth scenarios.
-This advantage stems from the robust denoising box structure, which enhances target location accuracy, and the use of central difference convolution to amplify target saliency.
-The foundation model SARATR-X, based on self-supervised learning with SAR images, performs well for fine-grained recognition and most of the EOC conditions, but the performance for scene and angle changes needs to be improved.
 Compared with the classification task, the detection task can effectively distinguish between target and background clutter, especially for target discovery in complex scenes such as woodlands.
@@
 However, the transfer effect is not obvious for the ship target class dataset in the sea scene.
+In recent years, the number of emerging satellite ship datasets has been much larger than that of other ground targets.
+This data can greatly compensate for the lack of high-resolution ground target samples in the existing SAR target datasets and advance the construction of an SAR ATR foundation model.
 4.3 Discussion
@@
 Robustness under different conditions.
-As Table IV and  V show, the variations in imaging conditions present a significant challenge for recognition.
+As Table V and  VI show, the variations in imaging conditions present a significant challenge for recognition.
 This variability in the domain distribution requires the extraction of stable invariant target features under different conditions.
@@
 Incorporation with multi-dimensional information.
-Using magnitude images alone loses many SAR image attributes, such as complex information that can effectively distinguish between targets and clutter in Table IV.
+Using magnitude images alone loses many SAR image attributes, such as complex information that can effectively distinguish between targets and clutter in Table V.
 Therefore, auxiliary information such as complex, polarization, and angle can be introduced in SAR ATR, but increasing the dimensionality of the information may increase the instability and complexity of the features.

### v3

@@
-NUDT4MSTAR: A Large Dataset and Benchmark Towards Remote Sensing Object Recognition in the Wild
+ATRNet-STAR: A Large Dataset and Benchmark Towards Remote Sensing Object Recognition in the Wild
 Abstract
-As an indispensable sensor for Remote sensing, Synthetic Aperture Radar (SAR) has a unique capability for all-day imaging.
-Nevertheless, in a data-driven era, the scarcity of large-scale datasets poses a significant bottleneck to advancing SAR automatic target recognition (ATR) technology.
-This paper introduces NUDT4MSTAR, a large-scale SAR dataset for remote sensing target recognition in the wild, including 40 vehicle target types and various imaging conditions across 5 realistic scenes.
-NUDT4MSTAR represents a significant leap forward in dataset scale, containing over 190,000 images—tenfold the size of its predecessors.
-We meticulously annotate each image with detailed target information and imaging conditions.
-Besides, data in both processed magnitude images and original complex formats are provided.
-Then, we construct a comprehensive benchmark consisting of 7 experiments with 15 recognition methods focusing on the stable and effective ATR issues.
-Besides, we conduct transfer learning experiments utilizing various models training on NUDT4MSTAR and apply them to three other target datasets, demonstrating its substantial potential for the broader field of ground objects ATR.
-Finally, we discuss this dataset’s application value and ATR’s significant challenges.
-To the best of our knowledge, this work marks the first-ever endeavor to create a large-scale dataset benchmark for fine-grained SAR recognition in the wild, featuring an extensive collection of exhaustively annotated vehicle images.
-We expect that the open source of NUDT4MSTAR will facilitate the development of SAR ATR and attract a wider community of researchers.
+The absence of publicly available, large-scale, high-quality datasets for Synthetic Aperture Radar Automatic Target Recognition (SAR ATR) has significantly hindered the application of rapidly advancing deep learning techniques, which hold huge potential to unlock new capabilities in this field.
+This is primarily because collecting large volumes of diverse target samples from SAR images is prohibitively expensive, largely due to privacy concerns, the characteristics of microwave radar imagery perception, and the need for specialized expertise in data annotation.
+Throughout the history of SAR ATR research, there have been only a number of small datasets, mainly including targets like ships, airplanes, buildings, etc.
+There is only one vehicle dataset MSTAR collected in the 1990s, which has been a valuable source for SAR ATR.
+To fill this gap, this paper introduces a large-scale, new dataset named ATRNet-STAR with 40 different vehicle categories collected under various realistic imaging conditions and scenes.
+It marks a substantial advancement in dataset scale and diversity, comprising over 190,000 well-annotated samples—10×10$\times$10 × larger than its predecessor, the famous MSTAR.
+Building such a large dataset is a challenging task, and the data collection scheme will be detailed.
+Secondly, we illustrate the value of ATRNet-STAR via extensively evaluating the performance of 15 representative methods with 7 different experimental settings on challenging classification and detection benchmarks derived from the dataset.
+Finally, based on our extensive experiments, we identify valuable insights for SAR ATR and discuss potential future research directions in this field.
+We hope that the scale, diversity, and benchmark of ATRNet-STAR can significantly facilitate the advancement of SAR ATR.
 Index Terms:
-Remote sensing, transformer, benchmark dataset, performance evaluation, synthetic aperture radar, automatic target recognition, image classification, object detection, foundation model, deep learning
+Remote sensing, transformer, datasets and benchmarks, performance evaluation, synthetic aperture radar, automatic target recognition, image classification, object detection, foundation model
 1 Introduction
-Figure 2: Motivation of our NUDT4MSTAR.
-Subfigure (a) depicts the most frequent keywords in 21,780 journal papers published in the remote sensing field (TGRS, J-STARS, GRSL, ISPRS P&R, and JAG) from 2020 to 2024.
-The size of each word is proportional to its frequency, highlighting that concepts such as synthetic aperture radar (SAR), image classification, and object detection [21, 10, 99] have garnered substantial attention.
+Figure 2: Motivation of our ATRNet-STAR.
+Subfigure (a) depicts the most frequent keywords in 21,780 journal papers published in the remote sensing field (TGRS, JSTARS, GRSL, ISPRS Journal, and JAG) from 2020 to 2024.
+The size of each word is proportional to its frequency, highlighting that concepts such as synthetic aperture radar (SAR), image classification, and object detection have garnered substantial attention.
 Subfigure (b) focuses on the number of publications related to SAR ATR over the past five years, a cross-area of the concepts highlighted in Subfigure (a).
@@
 However, the lack of large-scale datasets has significantly limited the growth of this research field in recent years.
-Remote sensing imagery interpretation is crucial for understanding our Earth comprehensively via diverse multi-modal sensors [64, 27, 111].
-Recently, the advent of the big data era has propelled the evolution of pre-trained foundation models, which have been widely applied across various remote sensing downstream tasks involving different multi-modal data [30, 107, 87, 26, 3, 58].
-Nevertheless, not all modal tasks in remote sensing possess a large number of target samples to exploit the potential of big data and foundation models.
-Synthetic aperture radar (SAR) [78, 36], an active imaging sensor, acquires the ground scattering characteristics by emitting the electromagnetic microwave.
-As shown in Fig 2 (a), it has emerged as an indispensable tool for information acquisition in Remote sensing and Earth observation [23, 83, 21], as its resolution remains unaffected by the observation distance or lighting conditions.
-SAR automatic target recognition (ATR), which autonomously detects and classifies objects of interest (e.g., , vehicles, ships, aircraft, and buildings) within imagery, is a critical increasing frontier in SAR interpretation, as illustrated in Fig.
-2 (b).
-It has various civil and military applications, such as urban management, disaster assessment, emergency rescue, and global surveillance [10, 54, 28, 108].
-Can humans perceive and recognize objects on the Earth beyond the limitations of optical vision using microwave vision [21, 36, 49]?
-Since the 1970s, researchers have been dedicated to developing a generalized stable SAR ATR system.
-However, recognizing objects within SAR images has encountered significant challenges in a big data era, especially the scarcity of large-scale target samples.
-Special characteristics.
-SAR imagery has special electromagnetic scattering attributions but is devoid of common features such as color and texture typically present in human vision.
-The scattering properties of targets are highly influenced by a multitude of factors, including the imaging angle, resolution, polarization, background, and the geometry structure and material of targets themselves [73, 44, 69].
-Additionally, SAR images suffer quality issues, such as speckle noise, geometric distortion, and image Defocus.
-These factors make it difficult to recognize objects in SAR, a unique recognition system independent of human vision.
-Insufficient samples.
-Acquiring SAR images encompassing many categories under diverse operating conditions presents a significant challenge, and annotating SAR images is considerably more costly and labor-intensive than RGB images.
-Three decades have elapsed since the advent of the first SAR classification dataset in the 1990s, the Moving and Stationary Target Acquisition and Recognition (MSTAR) [2].
-As shown in Fig.
-3, the largest publicly accessible SAR target classification dataset remains limited to a small set of fewer than 30,000 images.
-Besides, recent classification datasets (OpenSARShip [34] and FUSAR-Ship [33]) have severe sample imbalance, ranging from a few to thousands of samples per class.
-A substantial amount of work remains to be undertaken for SAR target datasets.
-Nonstandard benchmarks.
-A standardized dataset and experimental settings serve as the north star for developing advanced SAR ATR techniques.
-However, the mainstream MSTAR dataset in Fig.
-2 (b) has dozens of distinct experimental settings, and existing algorithms have nearly reached near-saturation performance on this ideal benchmark [44].
-Other classification datasets, such as OpenSARShip and FUSAR-Ship, also exhibit different experimental settings and dataset splitting.
-Moreover, the limiting access to non-open-source dataset settings and methods further intensifies this challenge.
-Researchers have struggled to construct many SAR target datasets to study generalized stable SAR ATR algorithms.
-Over the past few years, there has been a remarkable surge in remote sensing datasets across various modalities, including visible light [5, 56], SAR [57, 93], infrared [98], and multimodal [65].
-Recent SAR classification datasets in Table II have significantly improved data diversity.
-In particular, SAR detection datasets [57, 93] with 100,000 images have emerged.
-Yet, regarding SAR target recognition, no dataset has emerged that can completely surpass the MSTAR benchmark dataset, released in the 1990s, regarding the diversity of target types, imaging conditions, and data formats it offers.
-Notably, the lack of open-source availability, inadequate samples, long-tailed distributions, unabundant collection conditions, and independent experimental settings are still all issues that warrant attention.
-We suggest that a large-scale fine-grained SAR target dataset and benchmarks to investigate the electromagnetic scattering properties of targets across diverse conditions represent a critical and enduring endeavor.
+Synthetic Aperture Radar (SAR) imaging is capable of generating high resolution imagery irrespective of lighting conditions and weather [1, 2, 3, 4], and has become an indispensable tool for Earth Observation [5, 2, 6].
+SAR Remote Sensing (RS) data enables analysis and recognition of objects and scenes [2, 3, 7], which has become a valuable complement to other RS imagery [8, 9, 10].
+Consequently, as a fundamental and challenging field in RS image analysis, SAR automatic target recognition (ATR) [11, 12], which autonomously detects and classifies objects of interest (e.g., , vehicles, ships, aircraft, and buildings), has become an active research area for several decades (Fig.
+2 (a)).
+SAR ATR has a wide range of civilian and military applications, including military reconnaissance, urban management, disaster assessment, emergency rescue, and global surveillance [13, 14, 15, 16].
+Despite the remarkable achievement achieved over the past several decades in the field of SAR ATR, the accurate, robust, and efficient recognition of any target in an open world remains unresolved [11, 7].
+Need for ATRNet The advent of the big data era has propelled the evolution of RS pretraining foundation models [17, 18, 19, 20] where large-scale pre-training enables efficient cross-task adaptation with minimal finetuning.
+However, the scarcity of large-scale standardized datasets restricts the development of generalizable data-driven SAR ATR methods, compared to the success of foundation models in other RS sensors [21, 22, 23].
+The data sensitivity, acquisition costs, annotation difficulty, and complexity of SAR imaging hinder the establishment of open large-scale data ecosystems: More than 50% of SAR ATR studies still rely on the 1990s-era Moving and Stationary Target Acquisition and Recognition (MSTAR) [24] dataset (Fig.
+2 (b)).
+In addition, non-standardized evaluation protocols across existing benchmarks (e.g., MSTAR, OpenSARShip [25], and FUSAR-Ship [26]) impede objective algorithm comparison.
+Therefore, building a large SAR ATR dataset and benchmark is necessary to unlock new model capabilities in this field.
+We aim for ATRNet—a massive, diversified SAR target dataset benchmark for modeling and recognizing target characteristics.
+Need for ATRNet-STAR.
+Researchers have struggled to construct many SAR target datasets to study SAR ATR algorithms.
+Many SAR classification datasets in Table II have significantly improved data diversity.
+In particular, new SAR detection datasets [27, 28] have emerged with 100,000 images.
+However, our previous research [29, 30] on SAR foundation models revealed that collecting publicly available datasets yields fewer than 300,000 target samples, with severe sample imbalance: mainly ship detection datasets.
+As the inaugural phase of ATRNet, we focus on vehicle targets due to: (1) Recent datasets predominantly are targets with spaceborne SAR and constrained imaging geometries, whereas airborne platforms offer higher resolution (0.1-0.3m) and richer imaging conditions.
+(2) Vehicle recognition has driven SAR ATR research for three decades since the building of MSTAR, establishing mature research systems.
+(3) The 1990s-era MSTAR dataset, despite its seminal role, suffers from idealized acquisition conditions and saturated performance (near 99% accuracy), failing to reflect real-world complexities and support innovation in the 2020s.
 TABLE I: MSTAR v.s.
-NUDT4MSTAR.
-We can see the potential of the NUDT4MSTAR dataset as a new benchmark, surpassing MSTAR in terms of data diversity and challenges.
+ATRNet-STAR.
+We can see the potential of the ATRNet-STAR dataset as a new benchmark, surpassing MSTAR in terms of data diversity.
 AFRL: Air Force Research Laboratory.
@@
 # Samples: Number of samples.
-Dataset MSTAR NUDT4MSTAR
+Dataset MSTAR ATRNet-STAR
 Year 1995 2025
@@
 Platform Airborne Airborne
-[…]
+Resolution (m) 0.3 0.12 $\sim$ 0.15
 Band X X, Ku
@@
 Depression angle () 15, 17, 30, 45 15, 30, 45, 60
-[…]
+Azimuth angle () 0 $\sim$ 360 0 $\sim$ 360
 Annotation classification classification, detection
 # Samples 14,557 194,324
-As a first step towards building a large-scale SAR target dataset, we introduce NUDT4MSTAR [^1] The name NUDT4MSTAR comes from the abbreviation of the College of Electronic Science and Technology, the National University of Defense Technology, as Department IV.
-This dataset focuses on the finegrained single target recognition problem in SAR ATR, i.e., the image classification task, but we provide a horizontal bounding box label for the object detection task.
-NUDT4MSTAR: https://github.com/waterdisappear/NUDT4MSTAR., a fine-grained dataset for vehicle target recognition as a new breakthrough to the previous benchmark dataset MSTAR in Table.
+Our Objective and Solution.
+As a first step towards building a large-scale SAR target dataset, we introduce ATRNet-STAR [^1] STAR: Stationary Target Acquisition and Recognition.
+This dataset focuses on the fine-grained single target recognition problem in SAR ATR, i.e., the image classification task, but we provide a horizontal bounding box label for the object detection task in different condictions., a fine-grained dataset for vehicle target recognition as a new breakthrough to the previous benchmark dataset MSTAR in Table.
 I.
-In particular, NUDT4MSTAR collects 194,324 object images in the wild from 40 target types, 5 scenes, and various imaging conditions with detailed annotation.
+We completed program designation, data acquisition and processing, and benchmark construction with nearly two years of effort and considerable resources.
+In particular, ATRNet-STAR collects 194,324 object images from 40 target types, 5 scenes, and various imaging conditions with detailed annotation.
 It is the largest public SAR vehicle recognition dataset, even 10 times the size of any similar vehicle dataset previously available, as shown in Table.
 II.
-Then, to facilitate innovation and comparisons, we build a well-designed benchmark with 7 experimental settings and 15 recognition methods.
-Besides, the transfer learning results of various models trained on this dataset to other datasets show that our fine-grained recognition task benefits the feature learning of different ground targets.
+Its sufficient target samples can support multiple aspects of generation, detection, and classification research.
+To facilitate innovation and comparisons, we built a well-designed classification and detection benchmark, including 15 representative methods based on 7 experimental settings for robust ATR, few-shot, and transfer learning.
+Our results show that SAR ATR under complex conditions is still very difficult, and representative methods that performed well in previous datasets can not solve some of our difficult experimental conditions, such as image angle and scene variations.
+We also find the effectiveness of the pre-trained foundation model compared to other methods, which illustrates the importance of a large-scale dataset.
+Besides, the transfer learning results of various models pre-trained on this dataset to other datasets show that our dataset benefits the feature learning of different ground targets.
 Finally, We provide detailed analyses and discussions to explore important issues for future research.
-Our work shows the huge challenges and potential in SAR recognition and provides the benchmark for stable fine-grained ATR algorithms.
-The dataset and benchmark are available on our GitHub.
 We believe that our new dataset, with comprehensive imaging conditions and experiment benchmark, will further promote the promise and reproducible research in SAR ATR.
-The main contributions of this paper are three points:
-1.
-We build a large dataset for remote sensing object recognition in the Wild.
-NUDT4MSTAR includes 40 distinct vehicle target types and over 194,324 target samples across 5 realistic scenes and various imaging conditions.
-It provides a systematic study of the effects of different conditions on SAR target characteristics and promotes the development of new ATR research topics.
-2.
-We provide a comprehensive benchmark with 7 experimental settings and 15 recognition methods, covering both classification and detection recognition tasks.
-This benchmark offers a standardized experimental platform for researchers to compare the performance of different algorithms, facilitating innovation and development in SAR ATR technology.
-3.
-We demonstrate the NUDT4MSTAR dataset’s potential for foundation models.
-A transfer learning experiment based on various models pre-trained on this dataset shows that it is not only suitable for vehicle target recognition but also has the potential to support research in other related objects, promoting the application of SAR ATR in multiple situations.
+Our Contributions. The main contributions of this paper are listed below.
+Dataset and Benchmarks.
+To our best knowledge, the ATRNet-STAR is the largest dataset towards SAR vehicle recognition in the wild.
+Building ATRNet-STAR is a challenging task, and has taken nearly two years.
+We derive benchmark datasets from ATRNet-STAR to facilitate the performance evaluation of target classification and detection methods.
+Code library.
+To the best of our knowledge, we are the first to integrate publicly available and representative methods into a single code library to facilitate reproducible and large-scale performance evaluations in this field.
+Performance Evaluation.
+We review most major approaches and provide a comparative study of 15 representative methods with 7 experimental settings on classification and detection benchmarks derived from ATRNet-STAR, offering official baseline references when developing new approaches.
+Insights and Future Directions.
+By analyzing the quantitative results, we provide insightful discussion that can guide the development of SAR ATR algorithms and provide potential future research directions in this field.
+This work mainly focuses on the introduction, performance evaluation, and potential value of ATRNet-STAR.
+The full annotated dataset, code library, and reported results are available at https://github.com/waterdisappear/ATRNet-STAR after acceptance.
 The remainder of this paper is organized as follows.
@@
 Sec.
-3 introduces our dataset NUDT4MSTAR.
+3 introduces our dataset ATRNet-STAR.
 Secs.
@@
 2.1 Datasets for SAR target classification
-Datasets are the cornerstone of developing new techniques for SAR ATR, especially in a deep learning-based era.
-In the past decades, the MSTAR [2] dataset has played a key role in SAR target classification tasks.
+Datasets are the cornerstone of the development of new ATR techniques, especially in an era based on deep learning.
+In recent years, there has been notable progress in RS datasets across various modalities [31], including visible light [32, 33, 34, 35, 36], SAR [27, 28], infrared [37], and multimodal [38, 39].
+However, not all sensors possess sufficient target samples to fully leverage the potential of big data and foundation models.
+In the past decades, the MSTAR [24] dataset has played a pivotal role in SAR target classification tasks.
 However, current deep learning-based methods are saturated on this dataset benchmark setting.
-MSTAR’s insufficient data volume also suffers from data biases such as background correlation [44, 52].
-Therefore, many datasets have been built to advance the development of this field, especially in recent years.
-In order to provide a comprehensive analysis, we discuss SAR target classification datasets that focus on individual target characteristics from the 1990s to the 2020s, which are detailed listed in Table II and Fig.
+MSTAR’s insufficient samples also suffers from data biases such as background correlation [11, 40, 41].
+Therefore, many datasets been developed to promote the advancement of this field, especially in recent years.
+We analyze SAR target classification datasets that focus on individual target characteristics from the 1990s to the 2020s, which are detailed listed in Table II and Fig.
 3.
@@
 Dataset Year OA Source Band Pol. # Classes # Types Res. (m) Img. Size # Img. Description
-[…]
-[…]
+MSTAR [24] 1995 $\checkmark$ airborne X single 8 10 0.3 128$\sim$193 14,557 Most cited vehicle dataset
+QinetiQ [42] 2004 - airborne X quad - 9 0.3 100$\sim$150 4,006 Non-ideal background conditions
 CV Domes [43] 2010 $\checkmark$ simulation X quad 3 10 0.3 - - 3D simulation civil vehicle
@@
 SARSIM [45] 2017 $\checkmark$ simulation X - 7 14 0.1 139 21,168 Simulation CAD vehicle
-[…]
-[…]
+OpenSARShip [25] 2018 $\checkmark$ satellite C dual 16 - 2.7$\sim$22 9$\sim$445 26,679 Sentinel-1 ship dataset
+HR4S [46] 2019 - satellite C quad 21 - 3$\sim$25 - 1,962 GF-3 and RADARSAT-2 ship
 MGTD [47] 2019 - laboratory X single 1 2 0.01 128 1,728 Scaled vehicle models
 SAMPLE [48] 2019 $\checkmark$ simulation X single 7 10 0.3 128 2,690 Simulation and measured pairs
-[…]
+FUSAR-Ship [26] 2020 $\checkmark$ satellite C dual 98 - 1.1$\sim$1.7 512 5,243 High-resolution GF-3 ship
 IRIS-SAR [49] 2020 - simulation - - 8 355 - 512 63,900 Simulation CAD models
@@
 MATD [51] 2022 $\checkmark$ airborne Ku - 2 2 - 128 144 Multi-angle strimap SAR aircraft
-[…]
-[…]
+SAR-ACD [52] 2022 $\checkmark$ satellite C single 2 6 1 32$\sim$200 3,032 GF-3 aircraft dataset
+ATRNet-STAR 2024 $\checkmark$ airborne X, Ku quad 21 40 0.12 $\sim$ 0.15 128 194,324 A large-scale fine-grained dataset
 Figure 3: Timeline of SAR classification dataset.
 Compared to other datasets focusing on target slices in simple scenes, we provide diverse target samples from different scenes to study the more difficult issues, and it is a much larger SAR ATR dataset.
-Vehicle dataset.
-Vehicle datasets usually include various types of fine-grained vehicles and imaging conditions to study the robustness classification under different distributions in the training and test sets.
-We summarize the target classes and types of these SAR vehicle datasets in Fig.
-5.
+Vehicle datasets.
+Vehicle datasets typically include a wide variety of fine-grained vehicle types and diverse imaging conditions, designed to evaluate robust classification performance under varying training and test set.
+These SAR vehicle datasets are summarized in Fig.
+5, which provides an overview of the target categories.
 MSTAR was released by the Defense Advanced Research Projects Agency (DARPA) as the first public dataset for SAR target classification.
@@
 These two datasets are in the raw echo data format and can be used for SAR imaging research, but they are not as convenient for researchers to use as MSTAR.
-However, these two civilian vehicle datasets are not rich in target categories and acquisition conditions.
-SARSIM [66] and SAMPLE [47] are two simulation datasets used to study transfer learning.
+Besides, these two civilian vehicle datasets are not rich in target classes and acquisition conditions.
+SARSIM [45] and SAMPLE [48] are two simulation datasets designed to study transfer learning.
 In particular, SARSIM uses 14 CAD models, and SAMPLE has simulated and measured sample pairs based on partially MSTAR’s public vehicle.
@@
 MGTD [47] collects the training and test sets in two different laboratory backgrounds.
-Another issue is that the depression angles of MSTAR public images are mostly in 15° and 17°, and some categories of images are missing at other depression angles, such as 30° and 45°.
+Another issue is that the depression angles of MSTAR public images are mostly in 15 and 17, with some classes missing at other depression angles.
 Hence, NUAAminiSAR [50] collects images of military targets at different depression angles by circular SAR in a national defense park, including 8 military vehicles and an airplane.
-As shown above, no vehicle datasets are sufficiently rich in target types, scene settings, and sensor conditions.
-Ship dataset.
+As shown above, existing vehicle datasets lack sufficient diversity in target types, scene settings, and sensor conditions.
+Ship datasets.
 SAR ship datasets have emerged with the development of satellite SAR technology.
 The difference with the vehicle dataset is that the ship classification dataset usually uses a satellite platform, and the imaging angles are limited.
-Ship datasets can be constructed and labeled with global ship tracking systems, but label noise is a difficult problem.
+Ship datasets can be constructed and annotated using global ship tracking systems, but label noise remains a significant challenge.
 OpenSARShip [25], HR4S [46], and FUSAR-Ship [26] advance the development of ship classification methods.
@@
 Researchers construct various experimental settings and use some target categories for the classification based on these two datasets.
-Aircraft dataset.
-Aircraft datasets have emerged in recent years, presenting unique challenges in remote sensing.
+Aircraft datasets.
+Aircraft datasets have emerged in recent years, presenting unique challenges in RS.
 Due to their streamlined surfaces and less obvious scattering characteristics, aircraft are more difficult to classify than vehicle and ship targets.
 Multiangle Aircraft Target Dataset (MATD) [51] collects SAR images of two aircraft at different imaging angles with a drone strimap SAR equipment.
-SAR-ACD [79] contains 2 categories and 6 types of civil aircraft in 3 airports.
+SAR-ACD [52] contains 2 classes and 6 types of civil aircraft in 3 airports.
 Others.
@@
 Object detection dataset.
-SAR has many object detection datasets besides the classification datasets used for individual targets.
-These datasets are used for detecting multiple targets and have large image sizes.
-They contain more background interference and intensive targeting challenges.
-SAR object detection datasets similar to SARDet-100K [57] and FAIR-CSAR [93] with a volume of about 100,000 images are mainly satellite-based datasets, with a resolution of mostly greater than 1 m.
+Beyond the classification datasets focused on individual targets, numerous SAR object detection datasets are designed to detect multiple targets.
+These datasets typically feature large image sizes, significant background interference, and complex targeting challenges.
+Examples include SARDet-100K [27] and FAIR-CSAR [28], mainly satellite-based datasets containing approximately 100,000 images with resolutions exceeding 1 meter.
 This resolution makes it hard to detect and classify vehicle targets.
@@
 Therefore, our dataset provides a more fine-grained SAR vehicle characterization dataset under target class and imaging conditions.
-Due to the high collection and annotation cost, collecting and publishing target samples under different imaging conditions is very difficult and time-consuming.
-Even after close to thirty years since its inception, MSTAR remains the most used classification dataset.
+Collecting target samples under different imaging conditions remains a challenging and resource-intensive endeavor.
+After thirty years, MSTAR remains the most used classification dataset.
 The evolution of SAR ATR necessitates a new large-scale dataset and benchmark in the new era.
@@
     Richness.
-The dataset should include rich target categories, scenes, and imaging conditions to approximate the real world and reduce data bias.
-Besides, the dataset needs enough metadata, such as complex data, scene description, and imaging parameters.
-Researchers can propose new algorithms by combining metadata with SAR image characteristics.
+A comprehensive dataset should include diverse target categories, scenes, and imaging conditions to reflect real-world environments and reduce biases.
+It also needs to contain enough formats and metadata, such as complex data and imaging parameters, to help researchers in algorithm innovation.
   2. $\bullet$
@@
 (pseudo-color for better visualization of SAR magnitude images, and subfigures (b) are cropped for better visualization of target signature variations.)
-Therefore, we build and open-source a new large-scale dataset, NUDT4MSTAR, with various categories and imaging conditions.
+Therefore, we build and open-source a new large-scale dataset, ATRNet-STAR, with various categories and imaging conditions.
 This dataset is carefully labeled with extensive metadata and has an official benchmark based on 7 experimental settings and 15 comparison methods with 2 data formats.
 This dataset is our first step in building a new SAR target dataset system.
-In the future, we will expand the objects from vehicles to a wider category list to develop SAR ATR.
+In the future, we will expand the objects from vehicles to a wider category list to further promote SAR ATR research.
 2.2 Review of SAR Target Classification Algorithms
@@
 According to their different feature representations, we divided target recognition methods into four distinct paradigms: traditional features, deep learning, feature fusion, and classification with metadata.
-In an era increasingly dominated by data-driven algorithms, the availability of a rich, standardized, and usable dataset enriched with metadata is becoming increasingly pivotal.
+In an era increasingly dominated by data-driven algorithms, the availability of a rich, standardized, and usable dataset is becoming increasingly pivotal.
 Traditional features.
-Previous studies have proposed a plethora of valuable features, encompassing geometric structure features [38, 24, 13], electromagnetic-scattering features [70, 42, 102, 50], time-frequency features [77], local descriptors [15, 76], and sparse representations [32].
+Previous studies have proposed a wide range of valuable features, encompassing geometric structure features [54], electromagnetic-scattering features [55, 56, 57, 58], time-frequency features [59], local descriptors [7, 60], and sparse representations [61].
 These features are particularly tailored to the special coherent imaging mechanisms.
@@
 Deep learning algorithms excel at uncovering intricate patterns within datasets, particularly when ample data is available.
-Nowadays, these algorithms achieve a remarkable nearly 99% accuracy on MSTAR’s basic Standard Operating Condition (SOC) settings, but the Extended Operating Condition (EOC) with robustness and few-shot are more challenging problems.
-Convolutional Neural Network (CNN) [45, 63] is a common network in SAR target classification.
+Nowadays, these algorithms achieve remarkable accuracy, reaching nearly 99% on MSTAR’s basic Standard Operating Condition (SOC) settings.
+However, the Extended Operating Condition (EOC), which includes robustness and few-shot, remains a significant challenge.
+Convolutional Neural Network (CNN) [62, 63, 64] is a common network in SAR target classification.
 All-convolutional network (A-ConvNet) [65] was proposed to reduce overfitting and had high accuracy under many experiment settings but is not robust to noise.
-In recent years, Transformer [85, 16] has also been used in SAR target classification.
-After a decade of significant progress, SAR target recognition now incorporates many advanced deep learning techniques such as data augmentation [14, 46, 25], attention mechanism [71, 88], domain alignment [55], reinforcement Learning [40, 8], transfer learning [97], meta-learning [22], physical deep learning [11, 90, 37, 59], semi-supervised learning [106, 80], and self-supervised learning [101, 39, 53].
+In recent years, Transformer [66, 67, 68] emerged as a powerful alternative, offering new perspectives on feature extraction.
+After a decade of significant progress, SAR target recognition now incorporates many advanced deep learning techniques such as data augmentation [69, 70], attention mechanism [71, 72], domain alignment [53], reinforcement Learning [73, 74], transfer learning [75], meta-learning [76, 77, 78], physical deep learning [79, 80, 81], semi-supervised learning [82], and self-supervised learning [83, 29].
 The majority of research endeavors are based on the MSTAR dataset, and a subset of studies also create the different non-open source settings of OpenSARShip and FUSAR-Ship datasets.
@@
 [84] discussed the fusion of deep learning and sparse coding.
-FEC [103] used deep CNN and ASC features based on complex data with amplitude and phase.
+FEC [85] used deep CNN combined with ASC features derived from complex data, incorporating both amplitude and phase information.
 Li et al.
@@
 AaDRL [73] explored the SAR active recognition framework to fully exploit target characteristic variations at different azimuth angles based on the SAMPLE dataset.
-In addition, many multi-view methods [68, 81] were proposed with image sequences from different imaging angles.
+In addition, multi-view methods [93] were proposed with image sequences from different imaging angles.
 Since the target types are known, the researchers used 3D models and shape parameters as metadata for fine-grained tasks.
@@
 These configurations include image angle variation, target version/configure variation, noise corruption, occlusion, resolution variation, and background clutter variation [11, 53].
-In a similar vein, other SAR classification datasets, such as OpenSARShip and FUSAR-Ship, lack a standard experimental benchmark, and researchers also devise different bespoke settings that leverage subsets of the target class for classification [37, 59].
+In a similar vein, other SAR classification datasets, such as OpenSARShip and FUSAR-Ship, lack a standard experimental benchmark, and researchers also devise different bespoke settings that leverage subsets of the target class for classification [81].
 Evaluation metrics.
@@
 Accuracy quantifies the proportion of correctly classified samples relative to the total number of samples in the dataset.
-The confusion matrix describes the correspondence between the classification result and the ground truth.
-Within this matrix, each column corresponds to the predicted class labels, whereas each row represents the true class labels.
+The confusion matrix provides a detailed correspondence between the predicted and true class labels, with each column representing the predicted class and each row representing the true class.
 Besides, the Receiver Operating Characteristic (ROC) curve is a valuable tool for assessing used for binary classification and detection tasks such as suppressing background clutter and outlier rejection [65].
-Figure 5: Taxonomic systems of NUDT4MSTAR.
+Figure 5: Taxonomic systems of ATRNet-STAR.
 For civilian vehicles, the taxonomy is based on Chinese and European vehicle classification standards [96, 97], according to the vehicle’s purpose, structure, size, and mass.
 For military vehicles, we followed the MSTAR taxonomic system [98].
-(a) Taxonomic systems of NUDT4MSTAR.
-Our dataset is a comprehensive civilian vehicle map covering 4 main categories, 21 classes, and 40 types.
+(a) Taxonomic systems of ATRNet-STAR.
+Our dataset is a comprehensive civilian vehicle map covering 4 classes, 21 subclasses, and 40 types.
 We provide a detailed illustration of the histogram distribution for these 40 vehicle types, demonstrating the breadth and depth of our dataset.
@@
 In addition, information about vehicle types in the dataset is also available on the web.
-Compared to other datasets with long-tailed distributions, the MSTAR dataset has a more balanced target sample and diverse imaging conditions, which facilitates the establishment of experimental frameworks and the investigation of recognition challenges.
+Compared to other datasets with long-tailed distributions, the MSTAR dataset has a more balanced target sample and diverse imaging conditions.
 Therefore, MSTAR has accumulated a large number of algorithms and experimental settings.
-However, this dataset has not released all target samples, and the extant algorithms have achieved peak performance in prevalent settings, such as classification with ten target types, imaging angle variation, and version variation.
+However, this dataset has not released all target samples, and the extant algorithms have achieved peak performance in common popular settings [11], such as classification with ten target types, imaging angle variation, and version variation.
 The ideal acquisition conditions of the MSTAR dataset cannot show that the recognition problem is truly well solved.
-Scholars are endeavoring to address these limitations by proposing novel experimental settings and datasets.
+Scholars aim to address these limitations by proposing novel simulation experimental settings.
 However, as shown in Fig.
-4 (b), some simulation experimental settings based on measured data may not accurately reflect the real world.
-Besides, the sample diversity is difficult to ensure due to collection and labeling costs.
-Therefore, we perceive a constraint on advancing SAR ATR techniques due to data limitations.
-Constructing a new open-source dataset benchmark with diverse target categories and imaging conditions is the cornerstone of excellent and stable SAR ATR techniques.
-3 Building NUDT4MSTAR: Why, What, How
-In this section, we present NUDT4MSTAR from three perspectives: (1) Why we constructed this dataset in the 2020s; (2) What we did to ensure its diversity and standardization; (3) How the diversity and statistical analysis of this dataset to support SAR ATR.
-TABLE III: Statistics of NUDT4MSTAR.
+4 (b), some simulation experimental settings may not accurately reflect the real-world conditions.
+Besides, ensuring sample diversity is fraught with challenges due to the high data collection and labeling costs.
+These limitations collectively impose significant constraints on the advancement of SAR ATR techniques.
+Constructing a new open-source dataset benchmark with diverse target categories and imaging conditions is a challenging cornerstone of excellent and stable SAR ATR techniques.
+3 Building ATRNet-STAR: Why, What, How
+In this section, we present ATRNet-STAR from three perspectives: (1) Why we constructed this dataset in the 2020s; (2) What we did to ensure its diversity and standardization; (3) How the diversity and statistical analysis of this dataset to support SAR ATR.
+TABLE III: Statistics of ATRNet-STAR.
 We segment the targets with fixed-size slices and random offsets to investigate their characteristics and aim for robust recognition under different collection conditions.
@@
 # Img.: Number of images in ground and slant range coordinate systems.
-Scene # Types Platform Mode Band Res. (m) Pol. Dep. (°) Azi. (°) Img. Size # Img.
-[…]
-[…]
-[…]
-[…]
-[…]
+Scene # Types Platform Mode Band Res. (m) Pol. Dep. () Azi. () Img. Size # Img.
+City 40 airborne strimap X 0.12 $\sim$ 0.15 quad 15, 30, 45, 60 5 128 83,465
+Factory 40 airborne strimap X & Ku 0.12 $\sim$ 0.15 quad 15, 30, 45, 60 30 128 63,597
+Sandstone 40 airborne strimap X & Ku 0.12 $\sim$ 0.15 quad 15, 30, 45, 60 30 128 30,720
+Woodland 11 airborne strimap X & Ku 0.12 $\sim$ 0.15 quad 15, 30, 45, 60 30 128 8,094
+Bare soil 11 airborne strimap X & Ku 0.12 $\sim$ 0.15 quad 15, 30, 45, 60 30 128 8,448
 The emergence of new datasets and papers has led to rapid progress in various areas of SAR ATR, demonstrating the attention given to this issue in recent years.
@@
 We aim to achieve as much diversity as possible in target categories, scene settings, and sensor conditions.
-With this in mind, our dataset has huge potential to advance the field of SAR ATR significantly.
-Its values for SAR ATR are:
+Such multidimensional design establishes it as an instrumental resource for advancing SAR ATR research.
+Its principal values for the SAR ATR are fourfold:
   1. $\bullet$
@@
 Target.
-First of all, according to the Chinese and European vehicle classification standards [82, 19], we confirm the main target categories based on uses and structures of vehicles, including people carriers (car and bus), goods carriers (truck), and special purpose vehicles (special).
-Then, we subdivide each category into various target classes based on size and special structure and collect as many different types as possible.
-Finally, NUDT4MSTAR has 4 categories, 21 classes, and 40 types.
+First of all, according to the Chinese and European vehicle classification standards [96, 97], we confirm the main target classes based on uses and structures of vehicles, including people carriers (car and bus), goods carriers (truck), and special purpose vehicles (special).
+Then, we subdivide each class into various subclasses based on size and special structure and collect as many different types as possible.
+Finally, ATRNet-STAR has 4 classes, 21 subclasses, and 40 types.
 It covers most civilian vehicle categories and has the most richness on civil target classes compared to other SAR vehicle datasets.
-Figure 7: Demo images, classes, types, and sizes of forty vehicles in NUDT4MSTAR.
+Figure 7: Demo images, classes, types, and sizes of forty vehicles in ATRNet-STAR.
 Here are the SAR (pseudo-color for better visualization of SAR magnitude images) and corresponding RGB images.
@@
 Therefore, we classify various vehicle types based on their structure and size.
-Their length, width, and height measurements are listed in meters (m).
+Their size (length * width * height) measurements are listed in meters (m).
 Besides, it is worth noting that our SAR images are not in the ideal situation where the target is located right in the center of the demo image, just as in the MSTAR dataset, but has a random offset similar to the QinetiQ dataset.
@@
 We simultaneously gather data from satellites and airborne to study the platforms and resolutions.
-Table III shows the open-source data list.
+Table III presents the open-source data list.
 After a year and a half’s effort (from April 2023 to October 2024), we completed program designation, data acquisition and processing, and benchmark construction.
@@
 Annotation.
-We annotate the target types and positions with rectangular box labels according to the optical images and placement positions.
-Since we focus on individual target characteristics, the targets are placed at a certain distance from each other.
+Target categories and coordinates are annotated using rectangular box labels based on optical reference images and deployment records.
+Focusing on individual target signatures, all objects maintain specified separation distances during placement.
 Besides our vehicles, other vehicles in the scene are labeled as “other”.
@@
 […]
-[…]
+Data format.
+Due to the unique imaging mechanisms, processing steps, and image properties of SAR, we offer different levels of products to support the comprehensive study of the impact of imaging steps on recognition problems.
+We provide magnitude images in the ground range coordinate system with 8-bit and 32-bit uint formats.
+The magnitude images with 16-bit uint and complex data with 32-bit float in the slant range coordinate system are also available.
+In subsequent experimental benchmarks, we use 8-bit ground distance magnitude images and 32-bit slant distance complex data.
+We also perform other processing based on the slant distance data, such as phase gradient autofocus and polarization decomposition, but the current version of this dataset does not include the processed data after these steps.
+Raw echoes are also stored, but due to the large storage ($\sim$25TB), they will not be available for upload and download.
+Researchers can contact us to obtain raw echo if needed.
+We hope these products will help the researchers easily work with our dataset while providing more tractability to explore and discuss specific imaging mechanisms and image properties for SAR recognition applications.
 Metadata system.
@@
 3.4 Statistical Analysis
-We now discuss the properties of our dataset and provide some statistics compared with others.
+We present key characteristics and comparative statistics of our dataset relative to others.
 Class distribution.
-NUDT4MSTAR includes 4 categories, 21 classes and 40 types of vehicles with balanced samples, as shown in Fig.
+ATRNet-STAR includes 4 class, 21 subclasses and 40 types of vehicles with balanced samples, as shown in Fig.
 5.
@@
 Besides deploying corner reflectors for resolution measurements, we produce and place a reference target of multiple geometries as the “SLICY” target in the MSTAR dataset for scattering issues of a standard geometry.
-Other vehicle targets.
-Other vehicle targets in the scene are also labeled for the potential detection task in large remote images.
-However, this dataset mainly discusses the scattering properties of individual targets.
-These interfering vehicles can be used as part of a future rejection issue.
 Figure 9: Non-centered target slices. We apply the non-centered target setting following the QinetiQ dataset to increase detection difficulty.
@@
 Figure 10: Influence of different scenes on target characteristics in SAR images.
-Most current SAR target classification datasets are collected in simple scenes, e.g., (a) MSTAR is collected in a flat grass scene with a clean background and pseudo-correlation between target and background.
-While (b) our NUDT4MSTAR samples are collected under different positions in various scenes.
+Most current SAR target classification datasets are collected in simple scenes, e.g., , (a) MSTAR is collected in a flat grass scene with a clean background and pseudo-correlation between target and background.
+While (b) our ATRNet-STAR samples are collected under different positions in various scenes.
 In Fig.
@@
 Shadows from interfering objects in the former may obscure targets, while trees in the latter are likely to increase non-target scattering.
-Due to occlusion and layover are complex results of a combination of target, interference, and imaging geometry, we illustrate this problem with a single target demo.
+We illustrate this problem with a single target demo at the same angle across scenes, but occlusion and layover result from a combination of target, interference, and imaging geometry.
 These statistics are from an SUV Chang’an CSCS75 Plus (Vehicle 12) at different image angles in the 3 scenes’s ground range images.
 Occlusion in scenes.
-Previous vehicle datasets focus on target characteristics with simple scenes, and we collect samples from different scenes with the same target and angle.
+Whereas previous vehicle datasets prioritize target characteristics in ideal environments, we systematically acquire samples across diverse scenes.
 Fig.
@@
 The shadow of buildings and roadside trees in the factory may obscure targets and reduce target scattering, and the nearby trees in woodland have more obvious occlusion and increase non-target scattering.
-Figure 11: Target azimuth angle distribution of (a) MSTAR and (b) NUDT4MSTAR.
-Compared to the MSTAR dataset, which has depression angles mainly publicized at 17° and 15° with incomplete angles for most target classes, our NUDT4MSTAR dataset provides balance and comprehensive angles for all targets.
+Figure 11: Target azimuth angle distribution of (a) MSTAR and (b) ATRNet-STAR.
+Compared to the MSTAR dataset, which has depression angles mainly publicized at 17 and 15 with incomplete angles for most target classes, our ATRNet-STAR dataset provides balance and comprehensive angles for all targets.
 However, the sampling interval of target angles is sparser due to strip imaging.
@@
 11.
-However, the sampling interval of target azimuth angles is not equally dense in different scenes due to the limitation of stripmap imaging and cost problems.
+However, azimuth angle sampling density exhibits scene-dependent due to stripmap mode constraints and operational cost limitations.
 Figure 12: Ratio of strong and weak points in different bands and polarization.
@@
 In addition, 14 labelers and two inspectors perform the data annotation about four months.
+The entire project cost nearly two years.
 3.5 Dataset value
@@
 Robustness recognition.
-Our data has diverse acquisition conditions, and each image is labeled with specific information.
-Therefore, we can discuss in detail the robustness of the SAR recognition algorithms when the acquisition conditions (e.g., scene and sensor sensors) vary between the training and test sets.
+Our dataset encompasses diverse acquisition scenarios, with each image accompanied by exhaustive acquisition metadata.
+Therefore, we can discuss in detail the robustness of the SAR recognition algorithms when the acquisition conditions (e.g., , scene and sensor sensors) vary between the training and test sets.
 If we consider data from different acquisition conditions as domains, the training set can include multiple domains to learn domain invariant features, i.e., the domain generalization [100].
 Few-shot learning.
-In addition to acquisition condition variations, the problem of small samples due to high costs is also a concern issue.
-The few-shot setting can be combined with robust recognition, and our dataset’s large number of target types can enrich the task construction of meta-learning [31, 95, 94] settings in existing few-shot learning.
+Besides acquisition condition variation, the inherent challenge of limited training samples due to high SAR collection costs also presents critical research opportunities.
+The few-shot setting can be combined with robust recognition, and our dataset’s large number of target types can enrich the task construction of meta-learning [101, 102, 103, 104] settings in existing few-shot learning.
 Transfer learning.
@@
 Incremental learning.
-The diversity of our dataset can support the study of domain incremental learning and class incremental learning [84].
+The diversity of our dataset supports investigations into domain incremental learning and class incremental learning [107].
 The domain incremental learning can improve the algorithm’s robustness with a dynamic process, and class incremental learning can progressively increase the ability to recognize or reject new classes.
-This incremental capability is required for SAR ATR systems facing open environments.
+These incremental capacities are essential for SAR ATR in open environments.
 Physical deep learning.
 SAR images have unique properties, such as complex phase and polarization.
-Our dataset provides a variety of data formats that can support recognition studies based on these unique characteristics rather than relying only on quantified SAR magnitude images.
+Our dataset provides multi-format data to facilitate recognition studies leveraging these intrinsic attributes instead of only relying on quantized SAR magnitude images.
 In addition, detailed metadata can also provide more gains for SAR ATR.
+Generative Models.
+Beyond recognition tasks, our dataset enables controllable generation of target samples under diverse imaging conditions [2], as well as estimation of target parameters across varying acquisition scenarios.
 We encourage researchers to propose new experimental settings and research issues based on this dataset. Don’t hesitate to contact us if you get new ideas.
 4 Benchmarks
-We consider 7 experimental settings with 2 data formats as classification and detection benchmarks for the NUDT4MSTAR dataset in Table IV.
+We consider 7 experimental settings with 2 data formats as classification and detection benchmarks from this dataset in Table IV, named ATRBench.
 These experimental settings include 2 SOC settings sampled from similar distributions and 5 EOC settings with obvious distribution shifts.
-The data formats are magnitude images in the ground range coordinate system and complex images in the slant range coordinate system.
+The data formats consist of magnitude images in the ground range coordinate system and complex images in the slant range coordinate system.
 In conclusion, we evaluate our benchmarks under 7 experimental settings and 15 recognition methods with classification and detection tasks based on magnitude images and classification tasks based on complex images.
-These experiments provide valuable insights and huge potential into recognizing SAR target characteristics under different imaging conditions and fine-grained classes.
-Our dataset is sufficient to give a new challenging benchmark to advance SAR target recognition technology.
-We also welcome researchers to contact us to construct richer experimental benchmarks from original data and publish them with your clear insights.
+These experiments provide valuable insights and substantial potential into recognizing SAR target characteristics under different imaging conditions and fine-grained classes.
+Our dataset is sufficient to provide a new, challenging benchmark for advancing SAR target recognition technology.
+We also welcome researchers to contact us to construct richer experimental benchmarks from original data.
 4.1 Experimental settings
 Here, we describe the experimental settings in Table IV, including 2 SOC settings sampled from similar distributions and 5 EOC settings with obvious condition and distribution shifts.
-TABLE IV: SOC and EOC settings of NUDT4MSTAR.
+TABLE IV: SOC and EOC settings of ATRNet-STAR.
 The imaging conditions in SOC are similar, while EOC considers variations in a single imaging condition.
@@
 Setting Set # Types Scene Dep. Azi. Band Pol. # Img. (Ground) # Img. (Slant)
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+SOC-40 train 40 all 15, 30, 45, 60 0$\sim$360 X & Ku quad 68,091 67,780
+Test 40 all 15, 30, 45, 60 0$\sim$360 X & Ku quad 29,284 29,169
+SOC-50 train 50 all 15, 17, 30, 45, 60 0$\sim$360 X & Ku quad 18,071 18,071
+test 50 all 15, 30, 45, 60 0$\sim$360 X & Ku quad 17,603 17,613
+EOC-Scene train 40 simple 15, 30, 45, 60 0$\sim$360 X & Ku quad 19,584 19,584
+test 40 difficult 15, 30, 45, 60 0$\sim$360 X & Ku quad 77,791 77,365
+EOC-Depression train 40 all 15 0$\sim$360 X & Ku quad 24,361 22,206
+test 40 all 30, 45, 60 0$\sim$360 X & Ku quad 73,014 74,743
+EOC-Azimuth train 40 all 15, 30, 45, 60 0$\sim$60 X & Ku quad 18,636 18,592
+test 40 all 15, 30, 45, 60 60$\sim$360 X & Ku quad 78,739 78,357
+EOC-Band train 40 except city 15, 30, 45, 60 0$\sim$360 X quad 27,711 27,653
+test 40 except city 15, 30, 45, 60 0$\sim$360 Ku quad 27,763 27,732
+EOC-Polarization train 40 all 15, 30, 45, 60 0$\sim$360 X & Ku HH 24,361 24,246
+test 40 all 15, 30, 45, 60 0$\sim$360 X & Ku other 73,014 72,703
+SOC and EOC settings.
+SOC settings are those where the training and test sets have similar imaging conditions, and we have SOC-40 and SOC-50 settings.
+SOC-40 is created from a random data sample with a 7:3 training-to-test ratio.
+SOC-50 randomly selects data with a similar amount of MSTAR ten classes, and we combine our dataset with MSTAR.
+EOC settings are those where there is a significant domain shift between the training set and the test set, such as variation in imaging conditions and target state.
+For EOC-Scene, we use simple scenes (sandstone and bare soil) as the training set and complex scenes (urban, factory, and woodland) with occlusion as the test set.
+EOC-Depression is the training set with a 15 depression angle and the test set with 30, 45 and 60.
+EOC-Azimuth means that training has limited target azimuth angles within 0$\sim$60, and the test has other angles.
+Imaging angle and geometry variation can change target signatures and background clutter.
+EOC-band and EOC-polarization are designed to test the effect of different bands and change from HH-polarization to other polarizations on recognition, respectively.
+We do not consider some experimental settings that added simulation perturbations, such as random/Gaussian noise and simulation occlusions.
 Compared methods.
@@
 Here, we present the results of our classification, detection, and transfer learning tasks.
-TABLE V: Classification results.
-We use accuracy as a metric, i.e., the number of correctly classified samples in proportion to the total number of samples.
-Gray color methods use complex images, and magnitude image methods are sorted by network architecture, such as CNN and Transformer.
-Bolded text indicates the best result, while underlined text is the next best result.
+[…]
 Setting VGG16 [108] HDANet [53] ResNet18 [109] ResNet34 [109] ConvNeXt [64] ViT [67] HiViT [110] SARATR-X [30, 29] MS-CVNet [111] LDSF [90]
-SOC-40 88.8 89.1 90.6 91.7 96.0 76.4 86.8 96.4 80.9 88.2
-SOC-50 72.9 63.7 71.2 72.9 81.6 59.2 68.0 85.2 55.2 65.7
-EOC-Scene 21.6 33.6 16.1 18.0 16.5 12.9 15.8 19.5 26.1 29.7
-– (city) 22.7 33.7 18.1 18.9 17.9 11.6 13.9 20.4 26.1 32.7
-– (factory) 20.5 34.3 13.7 16.8 14.4 15.3 17.8 18.4 25.6 30.1
-– (woodland) 6.8 27.6 4.2 5.4 4.2 1.97 3.8 2.7 20.2 26.3
-EOC-Depression 33.2 32.9 33.9 37.2 43.1 30.4 31.4 39.9 21.8 28.3
-– (30) 52.4 54.0 52.2 55.9 63.2 43.7 47.3 58.1 39.2 46.5
-– (45) 33.5 32.7 34.2 38.3 45.1 31.0 32.0 41.2 19.2 25.9
-– (60) 14.0 12.3 15.4 17.4 21.3 16.7 15.0 20.5 8.1 12.7
-EOC-Azimuth 15.7 16.4 14.9 16.5 21.1 29.0 22.8 26.4 19.0 22.5
-– (60) 20.7 19.4 18.0 22.1 26.2 34.4 27.2 28.9 26.8 31.0
+SOC-40 $88.8$ $89.1$ $90.6$ $91.7$ $\underline{96.0}$ $76.4$ $86.8$ $\mathbf{96.4}$ $80.9$ $88.2$
+SOC-50 $72.9$ $63.7$ $71.2$ $72.9$ $\underline{81.6}$ $59.2$ $68.0$ $\mathbf{85.2}$ $55.2$ $65.7$
+EOC-Scene $21.6$ $\mathbf{33.6}$ $16.1$ $18.0$ $16.5$ $12.9$ $15.8$ $19.5$ $26.1$ $\underline{29.7}$
+– (city) $22.7$ $\mathbf{33.7}$ $18.1$ $18.9$ $17.9$ $11.6$ $13.9$ $20.4$ $26.1$ $\underline{32.7}$
+– (factory) $20.5$ $\mathbf{34.3}$ $13.7$ $16.8$ $14.4$ $15.3$ $17.8$ $18.4$ $25.6$ $\underline{30.1}$
+– (woodland) $6.8$ $\mathbf{27.6}$ $4.2$ $5.4$ $4.2$ $1.97$ $3.8$ $2.7$ $20.2$ $\underline{26.3}$
+EOC-Depression $33.2$ $32.9$ $33.9$ $37.2$ $\mathbf{43.1}$ $30.4$ $31.4$ $\underline{39.9}$ $21.8$ $28.3$
+– (30) $52.4$ $54.0$ $52.2$ $55.9$ $\mathbf{63.2}$ $43.7$ $47.3$ $\underline{58.1}$ $39.2$ $46.5$
+– (45) $33.5$ $32.7$ $34.2$ $38.3$ $\mathbf{45.1}$ $31.0$ $32.0$ $\underline{41.2}$ $19.2$ $25.9$
+– (60) $14.0$ $12.3$ $15.4$ $17.4$ $\mathbf{21.3}$ $16.7$ $15.0$ $\underline{20.5}$ $8.1$ $12.7$
+EOC-Azimuth 15.7 16.4 14.9 16.5 21.1 $\mathbf{29.0}$ 22.8 26.4 19.0 22.5
+– (60) 20.7 19.4 18.0 22.1 26.2 $\mathbf{34.4}$ 27.2 28.9 26.8 31.0
 – (120) 8.2 9.5 8.3 8.5 13.5 26.4 14.2 22.8 13.2 15.6
@@
 Figure 14: Transfer learning results.
-We report the 5-shot fine-tuning performance of transfer learning on three SAR target classification datasets after NUDT4MSTAR training.
+We report the 5-shot fine-tuning performance of transfer learning on three SAR target classification datasets after ATRNet-STAR training.
 The results show that our dataset has a good effect on ground target datasets such as vehicles and aircraft, but the effect is insignificant for sea targets.
@@
 4.3 Discussion
-Researchers have pursued realizing a stable and efficient SAR ATR recognition system, but previous vehicle benchmark datasets have not provided difficult challenges.
-Our results show that current state-of-the-art methods still face many robustness and small sample difficulties.
-Robust feature extraction under different imaging conditions and scenes is still a worthy research problem.
-It is also important to efficiently use a small number of labeled samples and mine complex and polarization information.
-Moreover, there is no single method that can handle different perturbation conditions, demonstrating the complexity of stable SAR ATR.
-Of course, building a large-scale dataset and advancing SAR ATR technology is a long-term process.
-In a data-driven approach dominant in the 2020s, a large-scale dataset will help to extract stabilizing features and be effective in many other datasets and tasks.
+Researchers have endeavored to realize a stable and efficient SAR ATR recognition system, but previous vehicle benchmark datasets have not provided sufficiently challenging conditions.
+Our results show that representative methods encounter significant challenges, such as imaging angle and scene variation, as well as few-shot settings.
+Robust feature extraction under different varying imaging conditions remains a critical research problem.
+It is also crucial to efficiently utilize a small number of labeled samples and mine complex and polarization information.
+Although the pre-trained foundation model SARATR-X is effective in most cases compared to other methods, no single method can effectively handle diverse perturbation conditions, highlighting the complexity of SAR ATR.
+In a data-driven approach dominant in the 2020s, our transfer learning results show that a large-scale dataset will help to extract stabilizing features and enhance performance across multiple datasets and tasks.
 Based on the results, we discuss several potential directions for the future development of ATR systems.
 Robustness under different conditions.
-As Table V and  VI show, the variations in imaging conditions present a significant challenge for recognition.
+As shown in Table V and  VI, the variations in imaging conditions present a significant challenge for recognition.
 This variability in the domain distribution requires the extraction of stable invariant target features under different conditions.
 Efficient utilization of samples.
-Due to the difficulty of acquiring and labeling SAR images, it is unlikely that the sample size will reach the same size as the visible light dataset.
+Given the challenges associated with acquiring and labeling SAR images, it is unlikely that the sample size will reach the same size as the visible light dataset.
 Therefore, it is important to improve learning efficiency by utilizing a small number of samples in Fig 13.
 Incorporation with multi-dimensional information.
-Using magnitude images alone loses many SAR image attributes, such as complex information that can effectively distinguish between targets and clutter in Table V.
-Therefore, auxiliary information such as complex, polarization, and angle can be introduced in SAR ATR, but increasing the dimensionality of the information may increase the instability and complexity of the features.
+Table V show that complex information can help distinguish between targets and clutter under scene variations.
+Auxiliary information such as complex and angle can be introduced in SAR ATR, but increasing the dimensionality of the information may increase the instability and complexity.
 Fast incremental learning capability.
-Because of the open and dynamic world, real ATR systems face constantly changing imaging conditions and target types.
+Because of the open real-world ATR systems face constantly changing imaging conditions and target types.
 They need to effectively recognize anomalies and update their capabilities in response to variations.
-This data provides a rich set of target types and imaging conditions that can be used to explore this possibility.
+This dataset provides a rich set of target types and imaging conditions that can be used to explore this possibility.
 Subsequently, we plan to construct a different resolution version and paired polarization data.
@@
 5 Conclusion
-A large-scale fine-grained SAR vehicle dataset and benchmark, named NUDT4MSTAR, has been established in this paper.
+A large-scale fine-grained SAR vehicle dataset and benchmark, named ATRNet-STAR, has been established in this paper.
 It contains various target types, scene variations, and imaging conditions and is 10 times the size of previous datasets of the same type.
-Our benchmark of 7 experiment settings and 15 methods result show that stable and efficient SAR ATR recognition method are still challengeable.
-It will advance diverse issue explorations and SAR ATR techniques with large-scale data.
+We provide a comprehensive benchmark of 7 experiment settings and 15 methods.
+It will facilitate diverse explorations of issues and promote SAR ATR techniques through large-scale data.
 In the future, we will continue to contribute to the SAR target datasets, and we welcome researchers to contact us to jointly promote the progress of the SAR ATR field.